### v1

# Past Movements-Guided Motion Representation Learning for Human Motion Prediction

Abstract
Human motion prediction based on 3D skeleton is a significant challenge in computer vision, primarily focusing on the effective representation of motion. In this paper, we propose a self-supervised learning framework designed to enhance motion representation. This framework consists of two stages: first, the network is pretrained through the self-reconstruction of past sequences, and the guided reconstruction of future sequences based on past movements. We design a velocity-based mask strategy to focus on the joints with large-scale moving. Subsequently, the pretrained network undergoes finetuning for specific tasks. Self-reconstruction, guided by patterns of past motion, substantially improves the model’s ability to represent the spatiotemporal relationships among joints but also captures the latent relationships between past and future sequences. This capability is crucial for motion prediction tasks that solely depend on historical motion data. By employing this straightforward yet effective training paradigm, our method outperforms existing state-of-the-art methods, reducing the average prediction errors by 8.8% across Human3.6M, 3DPW, and AMASS datasets. The code is available at https://github.com/JunyuShi02/PMG-MRL.
1 Introduction
3D skeleton-based motion prediction stands as a fundamental task in computer vision, with the objectives of generating consecutive and unseen future motion sequences based on past observations. This technique finds widespread application in domains such as human-computer interaction, autonomous driving, and motion analysis. The key to achieve accurate predictions of future motions lies in effectively modeling motion sequences, particularly the spatiotemporal representation of moving joints. However, due to the inherent properties of human kinematics, joints exhibit complex hierarchical rotational relationships. The motion trajectories of human joints also display a certain degree of randomness. Consequently, effectively capturing the spatiotemporal relationships between joints has perennially posed a challenging problem.
In the early stages, researchers utilized probabilistic models, such as Gaussian processes[38] and nonlinear Markov[19] models, to predict future motion changes based on historical motion sequences. However, these methods are computationally intensive and make strong assumptions about data distribution, which leads to models that are inadequate for real-world prediction scenarios. With the emergence of large-scale datasets[17, 37, 27] and advancements in hardware computing performance, motion prediction tasks started employing deep learning methods based on Recurrent Neural Networks (RNNs)[11, 16, 9, 14, 31]. RNN-based models effectively capture the dependencies between frames in a sequence, but the step-by-step generation approach results in the accumulation of prediction errors. The rise of Graph Neural Networks (GNNs) has brought feed-forward prediction paradigms into the motion prediction[26, 3, 10, 30, 34]. GNNs are well-suited for handling the non-Euclidean distribution properties of joints in spatial dimensions, enabling effective modeling of motion’s spatial information. Moreover, methods based on Transformers and Multi Layer Perceptrons (MLPs) have also achieved success in motion prediction tasks[1, 12, 2].
Overall, the mainstream paradigms for 3D skeleton-based motion prediction can be categorized into two main categories: 1) extracting features from historical motion sequences using an encoder and then generating predicted motions by a decoder[20, 11, 23] (as shown in Fig. 1(a)); 2) directly generating predicted motions using a feed-forward network[28, 26, 22] (as shown in Fig. 1(b)). However, the potential of self-supervised learning paradigms based on pretraining has not been sufficiently explored and applied in the field of motion prediction. This limits the ability of models to fully learn the spatiotemporal representations of motion sequences. AuxFormer[41] improves prediction performance by incorporating denoising tasks and masked feature prediction as auxiliary tasks during training (as shown in Fig. 1(c)). Nevertheless, parallel multi-task prediction encounters challenges such as interference with the main task (future motion prediction) and difficulties in balancing loss weights. Moreover, conducting masked feature prediction directly on the entire sequence lacks an exploration of the relationship between historical and future sequences.
In this paper, we propose a novel motion representation learning method for motion prediction tasks, which comprises two stages: pretraining and finetuning. In the pretraining stage, we apply masking to the joints of future motion sequences and utilize the complete past sequence features to guide the reconstruction of future sequences. This approach not only enhances the model’s reconstruction performance through past sequence features but also allows the model to learn the underlying relationship between historical and future sequences during the pretraining stage. Simultaneously, we also conduct self-reconstruction training based on masking on the past sequences. When applying the mask, we mask the joints with the largest motion changes between adjacent frames (i.e., the set of joints with the highest velocity). These joints serve as the central representation of the primary motion changes. Compared to random masking, this approach avoids learning motion patterns of joints with lower information content. For instance, in the ”Eating” action, motion changes are concentrated in the upper limb part, implying that we should focus on the representation learning of the upper limb part (including hand, head, and other joints), rather than the lower limb part with smaller changes.
Motion data contains both temporal and spatial dimensions. Given the outstanding performance of Transformers in temporal modeling tasks and spatial modeling tasks, as well as prior work on using Transformers for motion prediction tasks [41, 32, 1], we employ spatiotemporal Transformers to model motion sequences. Specifically, our method includes two components based on Transformers: the Past Motion Encoder (PME) and the Future Motion Predictor (FMP). The former comprises SpatioTemporal Attention (ST-Attention) modules, while the latter consists of ST-Attention modules and Past Motion Guided Attention (PMG-Attention) modules that fuse historical and future information. By pretraining the parameters of the PME and FMP, and finetuning them for the prediction task, our method significantly outperforms other state-of-the-art methods in both long-term and short-term predictions using a straightforward model architecture. It is worth noting that we observe a correlation between the performance improvement of our method and the increase in dataset size, which aligns with the characteristic of self-reconstruction learning requiring large-scale data.
The key contributions of our work can be summarized as follows:
Figure 1: (a) Prediction based on Encoder-Decoder paradigm[20] [23]. (b) Prediction based on Feed-forward paradigm[24] [28]. (c) Enhancing prediction based on auxiliary task[6] [41]. (d) Our Pretrain-Finetune prediction paradigm.
  • We propose a human motion representation learning method based on self-reconstruction pretraining paradigm for motion prediction tasks. This method helps the model learn the intrinsic spatiotemporal structure and features of motion sequences.
  • We use past motion features to guide the self-reconstruction of future sequences, which not only improves self-reconstruction performance but also helps the model understand the underlying relationship between past and future sequences in the motion prediction tasks.
  • We conduct extensive experiments and demonstrate that our method reduces the average prediction errors by 8.8% compared to state-of-the-art methods on the 3DPW, Human3.6M, and AMASS datasets.
2 Related Work
2.1 3D Skeleton-based Human Motion Prediction
RNN-based Methods
Early motion prediction methods employed probability model-based approaches[38, 19]. These methods performed poorly when dealing with complex motions. As RNNs can naturally process sequential data, researchers widely used RNNs and their variants for motion prediction tasks[11, 9, 14, 31]. EDR[9] maps motion sequences to a latent space using a nonlinear encoder and models temporal features using LSTM (Long Short-Term Memory Network). To avoid ambiguity when predicting long-term motions, VAE-LSTM[14] introduces the encoding of control signals into a variational inference framework trained to learn human motion manifolds. United[31] achieves smooth prediction by modeling first-order motion derivatives. To incorporate spatial information modeling, Structure-RNN[18] unfolds the spatiotemporal graph in the temporal dimension and decomposes it into a set of contributing factor components fed into RNN modeling. SkelNet[13], at each time frame, obtains local representations of human structure through special component layer branches.
Feed-forward Methods
The step-by-step modeling and generation approach of RNNs often leads to the issue of error accumulation. In recent times, many methods have adopted a feed-forward paradigm, utilizing GNN and Transformer models, to generate prediction sequences in a non-autoregressive manner[26, 3, 10, 30, 34]. These approaches leverage graph structures and attention computations that adapt to the non-Euclidean distribution properties of joints in space and time. LTD[28], HisRep[29], SGSN[22], and SPGSN[24] transform motion sequences to the frequency domain, modeling the trajectories of joints. DMGNN[20], MST-GNN[23], MSR-GCN[7], Sybio-GNN[21], and MMotion-Att[30] learn motion patterns at multiple levels by modeling human representations at different scales. DGMGNN[20], MANet[34], DD-GCN[40], and DSTD-GCN[10] model the dynamic characteristics of joints, adapting to the dynamic properties of motion. Transformer-based models utilize self-attention computations, which enhance the model’s capability to capture long-range dependencies between joints[1, 32, 41]. Among them, AuxFormer[41] achieves better results by incorporating denoising and reconstruction tasks to assist the prediction task. In addition, lightweight models based on MLP[2, 12] have also demonstrated success.
2.2 Self-supervised Pretraining
Self-supervised pretraining has brought about a revolution in the understanding and generation of language and images. This technique serves as one of the fundamental training methods of Large Language Models (LLMs), which is widely used in BERT[8], RoBERTa[25] and other models. For image processing tasks, recovering masked pixels/tokens is a common self-supervised pretraining method[15, 5]. MAE[15] proves that even when a substantial portion of image information is masked, the network can reconstruct a complete image based on a limited number of visible features. The self-learning paradigm of autoencoder in MAE is also widely used in multimodal[4], video processing[36], point cloud data processing[42] and other tasks.
3 Method
Figure 2: Detail of the model architecture. Our proposed model comprises Past Motion Encoder (PME) and Future Motion Predictor (FMP), where the former consists of $N\times$ SpatioTemporal Attention (ST-Attention), and the latter consists of $M$ ST-Attention and Past Motion Guided Attention (PMG-Attention). Here, the finetune stage is taken as an example to show the model structure (the pretrain stage only distinguishes the number of forward propagation of PME and incomplete input sequences).
3.1 Problem Background
The purpose of 3D skeleton-based human motion prediction is to generate the unseen and consecutive future motion based on past motion patterns. Assume a historical motion sequence consisting of $T$ frames in the Cartesian coordinate system, denoted as $\mathbf{X}^{P}=[\textbf{x}^{1},\textbf{x}^{2},...,\textbf{x}^{T}]\in\mathbb{R}^{T\times J\times K}$, where $J$ represents the number of joints and $K$ is the feature dimension (the initial value of $K$ is 3, representing the $x,y,z$ coordinates). Assume $\mathbf{X}^{F}=[\textbf{x}^{1},\textbf{x}^{2},...,\textbf{x}^{L}]\in\mathbb{R}^{L\times J\times K}$ is the future motion pose of $L$ frames. Our goal is to design and train a network $\mathcal{P}(\cdot)$ that takes $\mathbf{X}^{P}$ as input and $\mathbf{X}^{F}$ as output, i.e., $\mathbf{X}^{F}=\mathcal{P}(\mathbf{X}^{P},\pi_{\theta})$, where $\pi_{\theta}$ is the trainable parameters.
3.2 Model Architecture
As shown in Fig. 2, our proposed model consists of a Past Motion Encoder (PME) and a Future Motion Predictor (FMP). The PME aims to extract spatiotemporal features from the past motion sequence. The FMP aims to generate the future motion sequence based on the initial state (which consists of zero-valued inputs in motion prediction tasks) and past motion features.
Joint Embedder and Predict Head
Given an input sequence $\mathbf{X}$, we first use a single linear layer to embed each joint of each frame. To distinguish the spatiotemporal order information of the joints, we add trainable temporal and spatial position encoding to the motion embedding. This process can be represented as:
$\mathbf{H}=\mathbf{W}\mathbf{X}+\mathbf{b}+\mathbf{P}_{t}+\mathbf{P}_{s}$ (1)
where $\mathbf{W}\in\mathbb{R}^{K\times C},\mathbf{b}\in\mathbb{R}^{C}$ are the weights and biases of the linear layer, and $\mathbf{P}_{t},\mathbf{P}_{s}$ are the temporal and spatial position encoding, which are randomly initialized and trainable. $C$ denotes the number of feature channels. For the output of PME and FMP, a linear layer as the Predict Head is applied to obtain the 3D motion sequence.
Past Motion Encoder (PME)
We design PME to model the features of the past motion sequence. Human motion sequences contain two features: the change of joint position over time and the relationship between joints in space. We use the SpatioTemporal Attention (ST-Attention) to simultaneously model the spatiotemporal information of the joints. Specifically, PME contains $N$ cascaded ST-Attention modules, where each module first models the sequence dependencies in the temporal dimension and then models the joint relationships in the spatial dimension. Assuming the input to the $l^{th}$ layer of ST-Transformer is $\mathbf{H}^{l}$, the computation process can be expressed as:
$\mathbf{H}^{l+1}=\mathcal{F}_{attn}^{s}(\mathcal{F}_{attn}^{t}(\mathbf{H}^{l}))$ (2)
where $\mathcal{F}_{attn}^{s}(\cdot)$ / $\mathcal{F}_{attn}^{t}(\cdot)$ represents the spatial/temporal multi-head self-attention operation with the bottle neck.
Future Motion Predictor (FMP)
The purpose of FMP is to generate future motion sequences based on the initial state and the output from PME (i.e., past motion features). FMP comprises two modules: the ST-Attention module for modeling the sequence dependencies within itself and the Past Movement Guided Attention module (PMG-Attention). PMG-Attention applies cross-attention to incorporate past motion features into the current state, which can be mathematically represented as:
$\mathbf{H}^{l+1}=\mathcal{F}_{c-attn}^{s}(\mathcal{F}_{c-attn}^{t}(\mathbf{H}^{l},\mathbf{H}^{P}),\mathbf{H}^{P})$ (3)
where $\mathbf{H}^{P}$ is the past motion features output by PME. $\mathbf{H}^{l}$ is the input of $l^{th}$ PMG-Attention. $\mathcal{F}_{c-attn}^{s}(\cdot)$ / $\mathcal{F}_{c-attn}^{t}(\cdot)$ represents the spatial/temporal multi-head cross-attention calculation with the bottle neck. We provide more information about the ST-Attention and PMG-Attention in the supplementary materials (Sec. 7).
3.3 Past Movement Guided Pretraining
We apply the self-reconstruction pretraining paradigm to help the model learn the spatiotemporal representation of joints in motion sequences (as shown in the left part of Fig. 1(d)), which compromise three steps: 1) past motion reconstruction; 2) past motion encoding; 3) future motion reconstruction. We will introduce each step in the following.
Past Motion Reconstruction
Given the past motion sequence $\mathbf{X}^{P}\in\mathcal{R}^{T\times J\times K}$, we first calculate the motion velocity of each joint $\mathbf{V}=\mathbf{X}_{2:T}^{P}-\mathbf{X}_{1:T-1}^{P}\in\mathcal{R}^{(T-1)\times J\times K}$. We find the threshold $s$ in $\mathbf{V}$ such that the $r$of the values in $\mathbf{V}$ are greater than the threshold $s$. The mask of the past motion sequence $\mathbf{M}^{P}$ can be mathematically expressed as:
$\mathbf{M}^{P,(i,j)}=\left\{\begin{array}[]{ll}1&,\mathbf{V}^{(i,j)}\geq s\\ 0&,\mathbf{V}^{(i,j)}<s\end{array}\right.$ (4)
The masked past motion sequence can be computed by $\mathbf{X}_{masked}^{P}=\mathbf{X}^{P}\times\mathbf{M}^{P}$. Then, the joint embedders are applied to obtain the masked motion embeds $\mathbf{X}_{emb}^{P}$. Noting that we incorporate the trainable temporal and spatial position encoding into the motion embeds to learn the temporal order and spatial relationships of joints. Finally, we calculate the reconstructed motion sequence $\mathbf{X}_{rec}^{P}=\mathcal{F}_{PH}(\mathcal{F}_{PME}(\mathbf{X}_{emb}^{P}))$, where $\mathcal{F}_{PH}(\cdot)$ and $\mathcal{F}_{PME}(\cdot)$ are Predict Head and Position Motion Encoder.
Past Motion Encoding
This step aims to compute the features of complete past motion, which will be used to guide the future motion reconstruction and integrate the hidden relationship between the past sequence and the future sequence into the network. Given the complete input sequence $\mathbf{X}^{P}$, we first apply joint embedder to obtain the motion embeds $\mathbf{X}_{emb}^{P}$. Then, the past motion features can be computed by $\mathbf{H}^{P}=\mathcal{F}_{PME}(\mathbf{X}_{emb}^{P})$.
Future Motion Reconstruction
Given the future motion sequence $\mathbf{X}^{F}\in\mathcal{R}^{L\times J\times K}$, we first use the same method for masking and embedding the future motion sequence as for processing the past motion sequence. The masked future motion embeds, denoted as $\mathbf{X}_{emb}^{F}$ and the complete past motion features $\mathbf{H}^{P}$ are the input of FMP. In the PMG-Attention of FMP, past motion features guide the reconstruction of masked future motion through cross-attention mechanism. Finally, the reconstructed motion sequence can be expressed as: $\mathbf{X}_{rec}^{F}=\mathcal{F}_{PH}(\mathcal{F}_{FMP}(\mathbf{X}_{emb}^{F},\mathbf{H}^{P})))$.
3.4 Fine-tuning for Motion Prediction
In the finetune phase, we use the complete past motion sequence as input to the PME to get past motion features (as shown in the right part of Fig. 1(d)). In FMP, we use the initial state initialized with the value of 0 (but retain the position encoding) as input, and output the complete future sequence guided by the past motion features. Given the past motion sequence $\mathbf{X}^{P}$, this process can be represented as:
$\mathbf{H}^{P}$ $=\mathcal{F}_{JE}(\mathcal{F}_{PME}(\mathbf{X}^{P}))$ (5)
$\mathbf{H}^{F}$ $=\mathcal{F}_{JE}(\mathcal{F}_{FMP}(\mathbf{H}_{0}^{F},\mathbf{H}^{P}))$ (6)
$\mathbf{X}^{F}$ $=\mathcal{F}_{PH}(\mathbf{H}^{F})$ (7)
where $\mathbf{H}_{0}^{F}$ is the initial state initialized with the value of 0. $\mathcal{F}_{JE}$ represents the joint embedder.
3.5 Loss Function
Loss Function of Pretrain Stage
Pretrain stage performs past sequence reconstruction and future sequence reconstruction. Given the past and future reconstructed results $\hat{\mathbf{X}}_{rec}^{P}$ and $\hat{\mathbf{X}}_{rec}^{F}$, and the corresponding ground-truth (i.e. complete sequence) $\mathbf{X}_{rec}^{P}$ and $\mathbf{X}_{rec}^{F}$, the loss function can be represented as follow:
$\begin{split}\mathcal{L}_{pretrain}&=\frac{1}{TJ}\sum_{t=1}^{T}\sum_{j=1}^{J}||\mathbf{X}_{rec}^{P,(t,j)}-\hat{\mathbf{X}}_{rec}^{P,(t,j)}||_{F}^{2}\\ &+\frac{\alpha}{LJ}\sum_{l=1}^{L}\sum_{j=1}^{J}||\mathbf{X}_{rec}^{F,(l,j)}-\hat{\mathbf{X}}_{rec}^{F,(l,j)}||_{F}^{2}\end{split}$ (8)
where $\alpha$ is the hyperparameter to balance the loss weight.
Loss Function of finetune Stage
For finetune stage, given the predicted sequence $\hat{\mathbf{X}}$ and the ground-truth $\mathbf{X}$, the loss function can be defined as:
$\mathcal{L}_{finetune}=\frac{1}{LJ}\sum_{l=1}^{L}\sum_{j=1}^{J}||\mathbf{X}_{j}^{l}-\hat{\mathbf{X}}_{j}^{l}||_{F}^{2}$ (9)
Table 1: Comparisons of MPJPEs between our proposed method with other state-of-the-art methods for short-term prediction on Human3.6. We show the results of detailed actions such as walking and eating. The best results are highlighted in bold, and the second best results is underlined.
Motion Walking Eating Smoking Discussion
Milliseconds 80 160 320 400 80 160 320 400 80 160 320 400 80 160 320 400
DMGNN[20] 17.3 30.7 54.6 65.2 11.0 21.4 36.2 43.9 9.0 17.6 32.1 40.3 17.3 34.8 61.0 69.8
MSR[7] 12.2 22.7 38.6 45.2 8.4 17.1 33.0 40.4 8.0 16.3 31.3 38.2 12.0 26.8 57.1 69.7
Mixer[2] 12.0 23.0 42.4 51.3 7.8 16.5 34.3 43.0 7.4 15.2 30.3 37.6 10.7 25.0 56.9 71.1
PGBIG[26] 10.2 19.8 34.5 40.3 7.0 15.1 30.6 38.1 6.6 14.1 28.2 34.7 10.0 23.8 53.6 66.7
SPGSN[24] 10.1 19.4 34.8 41.5 7.1 14.8 30.5 37.9 6.7 13.8 28.0 34.6 10.4 23.8 53.6 67.1
siMLPe[12] 11.9 19.3 31.9 36.9 10.2 16.5 30.7 37.3 10.4 16.4 29.7 35.9 12.7 25.3 55.1 68.1
DSTD-GC[10] 11.1 22.4 38.8 45.2 7.0 15.5 31.7 39.2 6.6 14.8 29.8 36.7 10.0 24.4 54.5 67.4
AuxFormer[41] 8.9 16.9 30.1 36.1 6.4 14.0 28.8 35.9 5.7 11.4 22.1 27.9 8.6 18.8 38.8 49.2
CIST-GCN[33] 11.8 23.4 40.5 46.5 6.7 14.8 29.8 36.8 7.3 15.6 31.0 38.0 10.2 23.7 52.3 65.3
Ours 8.6 16.5 31.2 37.7 6.1 13.7 29.3 36.8 5.3 10.3 19.7 25.2 7.7 16.3 31.7 40.9
Motion Direction Greeting Phoning Posing
Milliseconds 80 160 320 400 80 160 320 400 80 160 320 400 80 160 320 400
DMGNN[20] 13.1 24.6 64.7 81.9 23.3 50.3 107.3 132.1 12.5 25.8 48.1 58.3 15.3 29.3 71.5 96.7
MSR[7] 8.6 19.7 43.3 53.8 16.5 37.0 77.3 93.4 10.1 20.7 41.5 51.3 12.8 29.4 67.0 85.0
Mixer[2] 7.5 18.4 43.9 55.3 16.4 37.1 80.2 98.2 9.1 19.5 42.4 53.6 11.5 27.0 64.9 83.5
PGBIG[26] 7.2 17.6 40.9 51.5 15.2 34.1 71.6 87.1 8.3 18.3 38.7 48.4 10.7 25.7 60.0 76.6
SPGSN[24] 7.4 17.2 40.0 50.3 14.6 32.6 70.6 86.4 8.7 18.3 38.7 48.5 10.7 25.3 59.9 76.5
siMLPe[12] 11.0 19.4 41.6 52.1 16.2 33.7 70.9 86.3 11.9 20.3 39.7 48.9 13.9 27.3 62.0 78.6
DSTD-GC[10] 6.9 17.4 40.1 51.7 14.3 33.5 72.2 87.3 8.5 19.2 40.3 49.9 10.1 25.4 60.6 77.3
AuxFormer[41] 6.8 17.0 40.3 51.6 13.5 31.3 69.2 85.4 7.9 17.3 37.4 47.2 8.8 19.1 39.2 51.0
CIST-GCN[33] 7.3 18.1 43.6 55.3 13.7 31.0 65.7 79.9 8.6 18.5 39.3 49.6 9.6 23.7 57.7 75.0
Ours 6.3 15.8 39.8 51.4 10.8 26.4 62.7 79.8 7.5 17.1 37.9 48.0 8.2 16.9 32.8 43.5
Table 2: Comparisons of MPJPEs between our proposed method with other state-of-the-art methods for short- and long- term prediction on Human3.6M. We show the average result of all actions. The best results are highlighted in bold.
Motion Average Param(M)
Milliseconds 80 160 320 400 560 1000 -
DMGNN[20] 17.0 33.6 65.9 79.7 103.0 137.2 4.82
MSR[7] 12.1 25.6 51.6 62.9 81.1 114.2 6.30
Mixer[2] 11.1 24.0 51.5 64.0 83.5 116.5 0.03
PGBIG[26] 10.3 22.7 47.4 58.5 76.9 110.3 1.80
SPGSN[24] 10.4 22.3 47.1 58.3 77.1 108.8 5.66
siMLPe[12] 13.4 23.7 47.4 58.0 77.9 111.7 0.14
DSTD-GC[10] 10.4 23.3 48.8 59.8 77.8 111.0 0.18
AuxFormer[41] 9.5 20.6 43.4 54.1 75.3 107.0 1.00
CIST-GCN[33] 10.5 23.2 47.9 59.0 77.2 110.3 0.35
GCNext [39] 9.3 21.5 45.5 56.4 74.7 108.7 -
Ours 8.8 19.3 41.1 51.8 70.0 101.5 1.66
4 Experiments
4.1 Dataset
Human3.6M
Human3.6M[17] is a large-scale, accurately annotated 3D human pose dataset that captures the motions of 11 professional actors in daily life scenarios using motion capture technology, totaling over 3.6 million pose instances. Following previous works[26, 24], only 22 joints are used, focusing on those that exhibit significant movement, and all sequences are down-sampled to 25 fps.
The 3D Poses in the Wild (3DPW)
3DPW[37] stands out as a pioneering dataset in motion related tasks, primarily due to its focus on capturing human motion in natural settings outside of the typical laboratory environment. This dataset provides a rich collection of video sequences captured at a frame rate of 30 frames per second, which includes diverse activities such as ’Walking’, ’Running’, and interacting with various objects.
The Archive of Motion Capture as Surface Shapes (AMASS)
AMASS[27] consolidates 15 existing motion capture datasets into a single, standardized framework. By converting these diverse datasets into a common representation, AMASS provides over 40 hours of meticulously annotated motion data. Following the previous works[34, 40], we discarded the detailed joints of the hand and sampled the sequence to 24fps.
Figure 3: Visualization results of short-term prediction (a) and reconstruction (b). The blue lines represent predicted/reconstructed motions, and the red lines represent ground-truths. We map the results to 2D space for display.
4.2 Experimental Settings
For the PME and the FMP, the number of modules (denoted as $N,M$) are both set to 3. The number of feature channels remains consistent at 128 throughout the network architecture. In the attention calculation, we employed 8 heads, with each head comprising 32 feature channels. During the pretraining phase, the loss weight (denoted as $\alpha$) is set to 1, indicating an equal contribution of loss from both past and future sequence reconstruction. We provide more implementation details in the supplementary materials (Sec. 8).
4.3 Comparison to State-of-the-art Methods
We compare our method with the recent state-of-the-art methods, including DMGNN (CVPR’20)[20, 30, 1, 7, 2, 26, 24, 12, 10, 41, 40, 33, 39]. To ensure fair comparisons, we re-evaluated some methods on the entire test set to align the testing standard using their official pretrained models, or retrain the model (Only for the cases where the official pre-training model is not provided) using their default hyperparameters on the Human3.6M.
Comparisons on Human3.6M
We use MPJPE (Mean Per Joint Position Error) as the evaluation metric, which is widely used in motion prediction tasks[26, 24, 41]. Tab. 1 shows the quantitative comparisons of short-term motion prediction on Human3.6M dataset, which aims to predict the poses within 400 milliseconds. Detailed actions are listed. As can be seen, our method outperforms the others on most actions, especially the ”Smoking”, ”Greeting”, and ”Posing”. Tab. 2 displays the average results of short - and long-term predictions for all actions on Human3.6M. Our approach effectively reduces prediction errors by 5.4%/10.2%/9.7%/8.2%/6.3%/6.6% at 80ms/160ms/320ms/400ms/560ms/1000ms.
Table 3: Comparisons of MPJPEs between our proposed method with other state-of-the-art methods for both short-term and long-term prediction on AMASS. The best results are highlighted in bold.
Milliseconds 80 160 320 400 560 720 880 1000
LTD-10-10[28] 10.3 19.3 36.6 44.6 61.5 75.9 86.2 91.2
LTD-10-25[28] 11.0 20.7 37.8 45.3 57.2 65.7 71.3 75.2
HisRep[29] 11.3 20.7 35.7 42.0 51.7 58.6 63.4 67.2
Mao W, et al.[30] 11.0 20.3 35.0 41.2 50.7 57.4 61.9 65.8
siMLPe[12] 10.8 19.6 34.3 40.5 50.5 57.3 62.4 65.7
DD-GCN[40] 10.4 19.1 33.6 39.8 49.3 56.5 61.3 64.6
Ours 8.6 15.3 27.7 33.2 42.0 48.3 53.0 56.0
Comparisons on AMASS
Tab. 3 shows the comparisons on AMASS dataset. Our method achieves a reduction in prediction errors of 17.3%/19.9%/17.6%/16.6%/14.8%/14.5%/13.5%/13.3% at 80ms/160ms/320ms/400ms/560ms/720ms/880ms/1000ms. We found that our method exhibits the most significant improvement on this dataset. This can be attributed to the larger data size of AMASS compared to Human3.6M and 3DPW. The self-reconstruction pretraining enables our model to learn a more comprehensive representation of motion on such a large dataset.
Table 4: Comparisons of MPJPEs between our proposed method with other state-of-the-art methods for short- and long-term prediction on 3DPW. The best results are highlighted in bold.
Average MPJPE
Milliseconds 100 200 400 600 800 1000
DMGNN[20] 17.80 37.11 70.38 94.12 109.67 123.93
HisRep[29] 15.88 35.14 66.82 93.55 107.63 114.75
MSR-GCN[7] 15.70 33.48 65.02 93.81 108.15 116.31
STSGCN[35] 18.32 37.79 67.51 92.75 106.65 112.22
PGBIG[26] 17.66 35.32 67.83 89.60 102.59 109.41
SPGSN[24] 15.39 32.91 64.54 91.62 103.98 111.05
AuxFormer[41] 14.21 30.04 58.50 89.45 100.78 107.45
Ours 13.85 29.11 57.06 87.35 97.93 103.37
Comparisons on 3DPW
The 3DPW dataset comprises data collected from outside the laboratory setting, which are widely used to verify the power of the model in real scenarios. As shown in Tab. 4, our approach performed best across all frames, showcasing its robustness in handling complex interactive scenarios.
Table 5: Comparison of Motion Classification Using Network-Output Features
w/o Network AuxFormer[41] Ours
Accuracy 25.6% 80.8% 95.0%
Representation Capability
To validate the representation learning capability of our proposed method, we first extract motion features from the outputs of the trained model (outputs from the penultimate layer). Subsequently, we train a simple linear classifier to classify these features. As shown in Tab. 5, our method significantly improves classification accuracy compared to methods that do not involve feature extraction via neural networks and AuxFormer[41].
4.4 Visualization of Prediction and Reconstruction
Fig. 3 (a) presents the visualization of prediction results. We conducted comparisons with PGBIG[26] and Mixer[2] on challenging samples. In the ’Posing’ and ’Discussion’ actions, our model successfully captures significant upward/downward movements of the arms. This can be attributed to the utilization of past motion guidance, which enables the model to comprehend the influence of previous motion patterns on future sequences. We also provide reconstruction results with a mask rate of 75% in Fig. 3 (b). Similar to the property of mask-based pretraining[15], even with a high mask rate of 75% for joints, the model can still accurately reconstruct the exact location of them.
5 Ablation Study
In this section, we explore the impact of components of our network, including the mask strategy, pretraining paradigm. We provide studies on model architecture in the supplementary materials. All experiments were carried out on the Human3.6M dataset.
Mask Strategy
Tab. 6 presents the results of two cases: A, without utilizing the velocity-based mask strategy (i.e. randomly mask 75% of the joints), and F, which represents the complete model. It can be seen that using the velocity-based mask reduced the average prediction errors by 2.3%. Tab. 6 also shows the comparisons of different mask rates (ranging from 0% to 75%), represented by B, C, D, F. The high mask rate facilitates the model to learn the space-time representation during the self-reconstruction process.
Effects of Pretraining Paradigm
We show the results of pretraining the model by denoising in Tab. 6 E. Compared with the self-reconstruction, denoising pretraining method lacks the exploration of the spatiotemporal relationship between joints.
Table 6: Impact of mask strategy and pretraining paradigm. $r$ is the mask rate. ’denoise’ means we conduct self-supervised pretraining by denoising.
ID 80ms 160ms 320ms 400ms
A random mask 9.0 19.7 42.4 52.8
B $r=0\%$ 9.6 21.9 46.2 56.8
C $r=25\%$ 9.3 20.2 42.9 53.7
D $r=50\%$ 9.1 20.0 42.3 52.6
E denoise 9.2 20.4 42.9 53.8
F complete ($r=75\%$) 8.8 19.3 41.1 51.8
6 Conclusion
In this paper, we propose a novel motion prediction paradigm, which consists two stages: 1) Past motion guided future motion self-reconstruction for learning the spatio-temporal representation of human joints (i.e. pretrain stage). 2) Train the network for the motion prediction task (i.e. finetune stage). Despite the simplicity of the model structure, our approach achieves the best results on three representative datasets and is highly extensible.
Supplementary Material
Table 7: Comparisons of MPJPEs between our proposed method with other state-of-the-art methods for short-term prediction on Human3.6. We show the results across all actions such as walking and eating. The best results are highlighted in bold, and the second best results is underlined
Motion Walking Eating Smoking Discussion Direction
Milliseconds 80 160 320 400 80 160 320 400 80 160 320 400 80 160 320 400 80 160 320 400
DMGNN[20] 17.3 30.7 54.6 65.2 11.0 21.4 36.2 43.9 9.0 17.6 32.1 40.3 17.3 34.8 61.0 69.8 13.1 24.6 64.7 81.9
MSR[7] 12.2 22.7 38.6 45.2 8.4 17.1 33.0 40.4 8.0 16.3 31.3 38.2 12.0 26.8 57.1 69.7 8.6 19.7 43.3 53.8
Mixer[2] 12.0 23.0 42.4 51.3 7.8 16.5 34.3 43.0 7.4 15.2 30.3 37.6 10.7 25.0 56.9 71.1 7.5 18.4 43.9 55.3
PGBIG[26] 10.2 19.8 34.5 40.3 7.0 15.1 30.6 38.1 6.6 14.1 28.2 34.7 10.0 23.8 53.6 66.7 7.2 17.6 40.9 51.5
SPGSN[24] 10.1 19.4 34.8 41.5 7.1 14.8 30.5 37.9 6.7 13.8 28.0 34.6 10.4 23.8 53.6 67.1 7.4 17.2 40.0 50.3
siMLPe[12] 11.9 19.3 31.9 36.9 10.2 16.5 30.7 37.3 10.4 16.4 29.7 35.9 12.7 25.3 55.1 68.1 11.0 19.4 41.6 52.1
DSTD-GC[10] 11.1 22.4 38.8 45.2 7.0 15.5 31.7 39.2 6.6 14.8 29.8 36.7 10.0 24.4 54.5 67.4 6.9 17.4 40.1 51.7
AuxFormer[41] 8.9 16.9 30.1 36.1 6.4 14.0 28.8 35.9 5.7 11.4 22.1 27.9 8.6 18.8 38.8 49.2 6.8 17.0 40.3 51.6
CIST-GCN[33] 11.8 23.4 40.5 46.5 6.7 14.8 29.8 36.8 7.3 15.6 31.0 38.0 10.2 23.7 52.3 65.3 7.3 18.1 43.6 55.3
Ours 8.6 16.5 31.2 37.7 6.1 13.7 29.3 36.8 5.3 10.3 19.7 25.2 7.7 16.3 31.7 40.9 6.3 15.8 39.8 51.4
Motion Greeting Phoning Posing Purchases Sitting
Milliseconds 80 160 320 400 80 160 320 400 80 160 320 400 80 160 320 400 80 160 320 400
DMGNN[20] 23.3 50.3 107.3 132.1 12.5 25.8 48.1 58.3 15.3 29.3 71.5 96.7 21.4 38.7 75.7 92.7 11.9 25.1 44.6 50.2
MSR[7] 16.5 37.0 77.3 93.4 10.1 20.7 41.5 51.3 12.8 29.4 67.0 85.0 14.8 32.4 66.1 79.6 10.5 22.0 46.3 57.8
Mixer[2] 16.4 37.1 80.2 98.2 9.1 19.5 42.4 53.6 11.5 27.0 64.9 83.5 13.4 30.7 65.6 80.0 9.8 20.9 45.6 57.3
PGBIG[26] 15.2 34.1 71.6 87.1 8.3 18.3 38.7 48.4 10.7 25.7 60.0 76.6 12.5 28.7 60.1 73.3 8.8 19.2 42.4 53.8
SPGSN[24] 14.6 32.6 70.6 86.4 8.7 18.3 38.7 48.5 10.7 25.3 59.9 76.5 12.7 28.6 61.0 74.4 9.3 19.4 42.2 53.6
siMLPe[12] 16.2 33.7 70.9 86.3 11.9 20.3 39.7 48.9 13.9 27.3 62.0 78.6 15.5 29.3 59.3 72.0 13.2 21.9 44.5 55.7
DSTD-GC[10] 14.3 33.5 72.2 87.3 8.5 19.2 40.3 49.9 10.1 25.4 60.6 77.3 12.7 29.7 62.3 75.8 8.8 19.3 42.9 54.3
AuxFormer[41] 13.5 31.3 69.2 85.4 7.9 17.3 37.4 47.2 8.8 19.1 39.2 51.0 11.9 28.0 61.8 76.3 8.7 19.0 42.1 53.3
CIST-GCN[33] 13.7 31.0 65.7 79.9 8.6 18.5 39.3 49.6 9.6 23.7 57.7 75.0 13.3 30.2 63.0 77.3 8.9 19.4 42.3 53.6
Ours 10.8 26.4 62.7 79.8 7.5 17.1 37.9 48.0 8.2 16.9 32.8 43.5 11.4 27.6 62.6 77.6 8.5 19.2 44.3 56.7
Motion Sitting Down Taking Photo Waiting Walking Dog Walking Together
Milliseconds 80 160 320 400 80 160 320 400 80 160 320 400 80 160 320 400 80 160 320 400
DMGNN[20] 15.0 32.9 77.1 90.3 13.6 29.0 46.0 58.8 12.2 24.2 59.6 77.5 47.1 93.3 160.1 171.2 14.3 26.7 50.1 63.2
MSR[7] 16.1 31.6 62.5 76.8 9.9 21.0 44.6 56.3 10.7 23.1 48.3 59.2 20.7 42.9 80.4 93.3 10.6 20.9 37.4 43.9
Mixer[2] 15.3 30.5 62.7 78.0 9.0 19.9 44.7 56.9 9.9 22.0 49.7 62.5 21.1 43.9 84.7 101.2 9.9 20.4 39.3 47.3
PGBIG[26] 13.9 27.9 57.4 71.5 8.4 18.9 42.0 53.3 8.9 20.1 43.6 54.3 18.8 39.3 73.7 86.4 8.7 18.6 34.4 41.0
SPGSN[24] 14.2 27.7 56.7 70.7 8.8 18.9 41.5 52.7 9.2 19.8 43.1 54.1 - - - - 8.9 18.2 33.8 40.9
siMLPe[12] 18.0 29.8 57.7 71.2 13.1 21.3 42.9 53.6 11.9 21.5 44.4 55.0 21.8 40.6 74.6 87.0 11.9 19.8 34.2 39.8
DSTD-GC[10] 14.1 28.0 57.3 71.2 8.4 18.8 42.0 53.5 8.7 20.2 44.3 55.3 19.6 41.8 77.6 90.2 9.1 19.8 36.3 42.7
AuxFormer[41] 13.5 27.6 57.7 72.2 8.2 18.4 41.5 53.0 8.2 18.5 41.2 52.2 17.1 36.5 70.4 83.0 7.8 15.9 30.2 37.0
CIST-GCN[33] 14.1 29.8 57.3 69.8 8.2 18.4 40.6 51.8 8.6 19.4 43.5 54.8 20.0 41.4 73.7 85.1 9.6 20.3 38.2 45.6
Ours 10.1 17.2 30.2 39.2 7.6 17.8 42.0 54.1 7.8 18.0 41.6 53.2 19.1 41.1 80.6 95.2 7.3 15.3 30.8 37.9
In this supplementary material, we provide more information that could not be included in the main manuscript because of space limit, including explanation of ST-Attention and PMG-Attention (in Sec. 7), implementation details (in Sec. 8), complete results for testing on Human3.6M dataset (in Sec. 9), ablation study on model architecture (in Sec. 10), and visualization of feature map (in Sec. 11).
7 Detailed Explanation of ST-Attention and PMG-Attention
In our model architecture for human motion prediction, both the Past Motion Encoder (PME) and the Future Motion Predictor (FMP) modules rely on attention mechanisms: SpatioTemporal Attention (ST-Attention) and Past Movement Guided Attention (PMG-Attention). These mechanisms are pivotal for capturing and integrating spatiotemporal dependencies in motion sequences.
ST-Attention
ST-Attention combines temporal and spatial attention mechanisms to model joint relationships over time effectively. This attention mechanism is utilized within PME to encode past motion features.
The temporal attention mechanism operates on the sequence dependencies over time in the joint positions:
$\mathbf{Q}^{t},\mathbf{K}^{t},\mathbf{V}^{t}=\mathcal{F}^{t}_{QKV}\left(\mathbf{H}^{l}\right)\$ (10)
$\mathbf{Z}^{t}=\operatorname{softmax}\left(\frac{\mathbf{Q}^{t}\left(\mathbf{K}^{t}\right)^{\mathrm{T}}}{\sqrt{d_{k}}}\right)\mathbf{V}^{t}$ (11)
where $\mathbf{H}^{l}$ represents the input at the $l^{th}$ layer, $\mathbf{Q}^{t},\mathbf{K}^{t}$, and $\mathbf{V}^{t}$ are queries, keys, and values, respectively, derived from $\mathbf{H}^{l}$, and $d_{k}$ is the dimension of keys.
The spatial attention mechanism focuses on modeling joint relationships in space:
$\mathbf{Q}^{s},\mathbf{K}^{s},\mathbf{V}^{s}=\mathcal{F}^{s}_{QKV}\left(\mathbf{Z}^{t}\right)$ (12)
$\mathbf{Z}^{s}=\operatorname{softmax}\left(\frac{\mathbf{Q}^{s}\left(\mathbf{K}^{s}\right)^{\mathrm{T}}}{\sqrt{d_{k}}}\right)\mathbf{V}^{s}$ (13)
where $\mathbf{Q}^{s},\mathbf{K}^{s}$, and $\mathbf{V}^{s}$ are queries, keys, and values for spatial attention, respectively, similarly derived from $\mathbf{Z}^{t}$. $\mathbf{Z}^{s}$ is the output of ST-Attention.
PMG-Attention
PMG-Attention is employed within the Future Motion Predictor (FMP) module to incorporate past motion features into the prediction of future motion sequences. Unlike ST-Attention which focuses on the past motion sequence itself, PMG-Attention applies cross-attention between the current state and the past motion features:
$\mathbf{Q}^{t},\mathbf{K}^{t},\mathbf{V}^{t}=\mathcal{F}^{t}_{QKV}\left(\mathbf{H}^{l}\right),\mathcal{F}^{t}_{QKV}\left(\mathbf{H}^{P}\right)$ (14)
$\mathbf{Z}^{t}=\operatorname{softmax}\left(\frac{\mathbf{Q}^{t}\left(\mathbf{K}^{t}\right)^{\mathrm{T}}}{\sqrt{d_{k}}}\right)\mathbf{V}^{t}$ (15)
$\mathbf{Q}^{s},\mathbf{K}^{s},\mathbf{V}^{s}=\mathcal{F}^{s}_{QKV}\left(\mathbf{Z}^{t}\right),\mathcal{F}^{s}_{QKV}\left(\mathbf{H}^{P}\right)$ (16)
$\mathbf{Z}^{s}=\operatorname{softmax}\left(\frac{\mathbf{Q}^{s}\left(\mathbf{K}^{s}\right)^{\mathrm{T}}}{\sqrt{d_{k}}}\right)\mathbf{V}^{s}$ (17)
where $\mathbf{Q}^{t}$ and $\mathbf{Q}^{s}$ are obtained from the hidden state of future motions, and the $\mathbf{K}^{t},\mathbf{V}^{t},\mathbf{K}^{s},\mathbf{V}^{s}$ are obtained from the features of past motion.
8 Implementation Details
We deployed our network using an NVIDIA RTX4090 GPU (with PyTorch 2.0). For training, we set the batch size to 24 and utilized the Adam optimizer with an initial learning rate of 0.0005. To enhance convergence, we implemented a cosine annealing learning rate schedule. These training parameters were consistently applied across all datasets in our experiments. We ensured that the GPU memory usage remained below 6GB, enabling our network to be seamlessly deployed on various hardware platforms with ease.
9 Complete Results on Human3.6M
We provide the full results of comparisons on the Human3.6M dataset in Tab. 7 as a supplement to Tab. 1. Despite the simplicity of our network structure, we achieved the best results on most actions.
10 Ablation Study on Model Architecture
Table 8: Ablation study of model architecture
ID 80ms 160ms 320ms 400ms
A add 9.4 20.1 42.7 53.5
B connect 9.2 20.1 42.7 53.7
C parallel attn. 9.1 20.1 42.3 52.9
D complete 8.8 19.3 41.1 51.8
Way of Fusing Past Motion and Future Motion
In the default setting, we use cross-attention to integrate features of past and future motions. We experimented with other fusion methods, with results shown in Tab. 8: A) Direct addition of features from past and future actions; B) Concatenation of both in the feature dimension, followed by a Linear layer to reduce the number of channels back to the original value. We observed that cross-attention is the best choice because it explicitly computes the temporal and spatial interactions between the past and future motion.
Table 9: Impact of model size. $d$ is the feature dimension. $l$ is the number of layers.
ID 80ms 160ms 320ms 400ms Param
E $d=64$ 9.3 20.2 43.3 54.4 0.67M
F $d=128$ 9.1 19.8 42.0 52.9 1.00M
G $l=2$ 9.2 20.2 42.6 53.3 1.16M
H $l=4$ 8.8 19.3 41.6 52.3 2.15M
D complete 8.8 19.3 41.1 51.8 1.66M
Figure 4: Visualization of attention feature maps.
Spatiotemporal Attention
In the default setting, we follow the AuxFormer[41] of computing temporal and spatial attention sequentially. Results of parallel computation of temporal and spatial attention are presented in Tab. 8 C, showing worse performance.
Model Size
In Tab. 9, we show the impact of different model sizes (including feature dimension $d$ in attention layer and number of layers $l$) on parameter count and results. Although reducing the model size can decrease the number of parameters, it can also lead to decreased performance. We observe that the default setting ($d$=256, $l$=3) represents the optimal balance between parameter count and performance.
11 Visualization of Attention Feature Map
For the example of the Discussion action, we present the averaged attention feature maps over 512 samples. As shown in Fig. 4, sub figure (a) and (b) display the temporal self-attention and spatial self-attention feature maps from different layers in PME, sub figure (c) and (d) show the temporal cross-attention and spatial cross-attention feature maps from different layers in FMP.
In temporal cross-attention, the model’s first and third layers tend to allocate attention to the later frames, as these frames have a greater influence on future sequences. In the second layer, the model distributes attention to both later and earlier frames, highlighting the potential impact of historical sequences on future actions. In spatial cross-attention, the model identifies significant correlations between the movement patterns of the ’thorax’, ’head’, ’left elbow’, and ’right elbow’ joints and the prediction of future actions.

### v2

@@
-Past Movements-Guided Motion Representation Learning for Human Motion Prediction
+MoReFun: Past-Movement Guided Motion Representation Learning for Future Motion Prediction and Understanding
 Abstract
-Human motion prediction based on 3D skeleton is a significant challenge in computer vision, primarily focusing on the effective representation of motion.
-In this paper, we propose a self-supervised learning framework designed to enhance motion representation.
-This framework consists of two stages: first, the network is pretrained through the self-reconstruction of past sequences, and the guided reconstruction of future sequences based on past movements.
-We design a velocity-based mask strategy to focus on the joints with large-scale moving.
-Subsequently, the pretrained network undergoes finetuning for specific tasks.
-Self-reconstruction, guided by patterns of past motion, substantially improves the model’s ability to represent the spatiotemporal relationships among joints but also captures the latent relationships between past and future sequences.
-This capability is crucial for motion prediction tasks that solely depend on historical motion data.
-By employing this straightforward yet effective training paradigm, our method outperforms existing state-of-the-art methods, reducing the average prediction errors by 8.8% across Human3.6M, 3DPW, and AMASS datasets.
-The code is available at https://github.com/JunyuShi02/PMG-MRL.
+3D human motion prediction aims to generate coherent future motions from observed sequences, yet existing end-to-end regression frameworks often fail to capture complex dynamics and tend to produce temporally inconsistent or static predictions—a limitation rooted in representation shortcutting, where models rely on superficial cues rather than learning meaningful motion structure.
+We propose a two-stage self-supervised framework that decouples representation learning from prediction.
+In the pretraining stage, the model performs unified past–future self-reconstruction, reconstructing the past sequence while recovering masked joints in the future sequence under full historical guidance.
+A velocity-based masking strategy selects highly dynamic joints, forcing the model to focus on informative motion components and internalize the statistical dependencies between past and future states without regression interference.
+In the fine-tuning stage, the pretrained model predicts the entire future sequence, now treated as fully masked, and is further equipped with a lightweight future-text prediction head for joint optimization of low-level motion prediction and high-level motion understanding.
+Experiments on Human3.6M, 3DPW, and AMASS show that our method reduces average prediction errors by 8.8% over state-of-the-art methods while achieving competitive future-motion understanding performance compared to LLM-based models.
+Code is available at: https://github.com/JunyuShi02/MoReFun
 1 Introduction
-3D skeleton-based motion prediction stands as a fundamental task in computer vision, with the objectives of generating consecutive and unseen future motion sequences based on past observations.
-This technique finds widespread application in domains such as human-computer interaction, autonomous driving, and motion analysis.
-The key to achieve accurate predictions of future motions lies in effectively modeling motion sequences, particularly the spatiotemporal representation of moving joints.
-However, due to the inherent properties of human kinematics, joints exhibit complex hierarchical rotational relationships.
-The motion trajectories of human joints also display a certain degree of randomness.
-Consequently, effectively capturing the spatiotemporal relationships between joints has perennially posed a challenging problem.
-In the early stages, researchers utilized probabilistic models, such as Gaussian processes[38] and nonlinear Markov[19] models, to predict future motion changes based on historical motion sequences.
-However, these methods are computationally intensive and make strong assumptions about data distribution, which leads to models that are inadequate for real-world prediction scenarios.
-With the emergence of large-scale datasets[17, 37, 27] and advancements in hardware computing performance, motion prediction tasks started employing deep learning methods based on Recurrent Neural Networks (RNNs)[11, 16, 9, 14, 31].
-RNN-based models effectively capture the dependencies between frames in a sequence, but the step-by-step generation approach results in the accumulation of prediction errors.
-The rise of Graph Neural Networks (GNNs) has brought feed-forward prediction paradigms into the motion prediction[26, 3, 10, 30, 34].
-GNNs are well-suited for handling the non-Euclidean distribution properties of joints in spatial dimensions, enabling effective modeling of motion’s spatial information.
-Moreover, methods based on Transformers and Multi Layer Perceptrons (MLPs) have also achieved success in motion prediction tasks[1, 12, 2].
-Overall, the mainstream paradigms for 3D skeleton-based motion prediction can be categorized into two main categories: 1) extracting features from historical motion sequences using an encoder and then generating predicted motions by a decoder[20, 11, 23] (as shown in Fig.
-1(a)); 2) directly generating predicted motions using a feed-forward network[28, 26, 22] (as shown in Fig.
-1(b)).
-However, the potential of self-supervised learning paradigms based on pretraining has not been sufficiently explored and applied in the field of motion prediction.
-This limits the ability of models to fully learn the spatiotemporal representations of motion sequences.
-AuxFormer[41] improves prediction performance by incorporating denoising tasks and masked feature prediction as auxiliary tasks during training (as shown in Fig.
-1(c)).
-Nevertheless, parallel multi-task prediction encounters challenges such as interference with the main task (future motion prediction) and difficulties in balancing loss weights.
-Moreover, conducting masked feature prediction directly on the entire sequence lacks an exploration of the relationship between historical and future sequences.
-In this paper, we propose a novel motion representation learning method for motion prediction tasks, which comprises two stages: pretraining and finetuning.
-In the pretraining stage, we apply masking to the joints of future motion sequences and utilize the complete past sequence features to guide the reconstruction of future sequences.
-This approach not only enhances the model’s reconstruction performance through past sequence features but also allows the model to learn the underlying relationship between historical and future sequences during the pretraining stage.
-Simultaneously, we also conduct self-reconstruction training based on masking on the past sequences.
-When applying the mask, we mask the joints with the largest motion changes between adjacent frames (i.e., the set of joints with the highest velocity).
-These joints serve as the central representation of the primary motion changes.
-Compared to random masking, this approach avoids learning motion patterns of joints with lower information content.
-For instance, in the ”Eating” action, motion changes are concentrated in the upper limb part, implying that we should focus on the representation learning of the upper limb part (including hand, head, and other joints), rather than the lower limb part with smaller changes.
-Motion data contains both temporal and spatial dimensions.
-Given the outstanding performance of Transformers in temporal modeling tasks and spatial modeling tasks, as well as prior work on using Transformers for motion prediction tasks [41, 32, 1], we employ spatiotemporal Transformers to model motion sequences.
-Specifically, our method includes two components based on Transformers: the Past Motion Encoder (PME) and the Future Motion Predictor (FMP).
-The former comprises SpatioTemporal Attention (ST-Attention) modules, while the latter consists of ST-Attention modules and Past Motion Guided Attention (PMG-Attention) modules that fuse historical and future information.
-By pretraining the parameters of the PME and FMP, and finetuning them for the prediction task, our method significantly outperforms other state-of-the-art methods in both long-term and short-term predictions using a straightforward model architecture.
-It is worth noting that we observe a correlation between the performance improvement of our method and the increase in dataset size, which aligns with the characteristic of self-reconstruction learning requiring large-scale data.
-The key contributions of our work can be summarized as follows:
-Figure 1: (a) Prediction based on Encoder-Decoder paradigm[20] [23].
-(b) Prediction based on Feed-forward paradigm[24] [28].
-(c) Enhancing prediction based on auxiliary task[6] [41].
-(d) Our Pretrain-Finetune prediction paradigm.
-  • We propose a human motion representation learning method based on self-reconstruction pretraining paradigm for motion prediction tasks.
-This method helps the model learn the intrinsic spatiotemporal structure and features of motion sequences.
-  • We use past motion features to guide the self-reconstruction of future sequences, which not only improves self-reconstruction performance but also helps the model understand the underlying relationship between past and future sequences in the motion prediction tasks.
+Figure 1: Comparison of complex motion prediction results.
+While previous methods struggle with complex movements, producing poor and unnatural static poses, our method demonstrates improved accuracy by addressing the inherent weakness in representation capability, yielding a natural motion prediction.
+3D human pose prediction, which aims to generate continuous and natural future motions based on historical observed sequences, is a pivotal technology for applications such as human-computer interaction and autonomous driving.
+Despite considerable strides made by recent studies, particularly those based on GNNs [ma2022progressively, cai2020learning, fu2023learning, mao2021multi, shi2023multi] and Transformers [aksan2021spatio, guo2023back, wei2025alien], existing state-of-the-art (SOTA) models exhibit fundamental failures in accurately predicting complex motions.
+Their predicted motions often rapidly degenerate into unnatural static poses or lose temporal coherence, as shown in the Fig.
+1.
+This failure is not attributable to specific models (e.g.
+Transformers) but rather stems from a fundamental flaw in the prevailing end-to-end regression paradigm used for motion prediction.
+Current methods frame this task as a direct regression problem.
+This approach compels the model to simultaneously learn complex spatio-temporal representations and regress a highly uncertain future within a single training stage.
+This results in a phenomenon we term “Representation Shortcutting” where the model learns only superficial features sufficient to minimize the regression loss on the training data, while failing to capture the underlying physical principles and joint constraints governing motion.
+To improve the prediction performance, researchers have pursued two main paths.
+The first path involves refining the regression paradigm.
+For instance, feed-forward architectures [ma2022progressively, xu2023eqmotion, sun2024moml] have been used to mitigate temporal error accumulation, while other methods model low-frequency DCT signals to reduce motion jitter [wang2024gcnext, li2021skeleton].
+Moreover, the system latency [sun2025lal] and unexpected perturbation [yue2024human] are also under consideration.
+However, these methods remain constrained by the “Representation Shortcutting” problem.
+The second path explores auxiliary learning, such as using parallel denoising tasks to aid prediction [xu2023auxiliary].
+Yet, this parallel multi-task paradigm has critical flaws: (1) The auxiliary and main (motion prediction) tasks may suffer from mutual interference; and (2) It fails to explicitly model the fundamental dependency of future motion on past observations, which is central to accurate motion prediction.
+In this work, we decouple representation learning from the prediction task into two distinct stages.
+In the first stage, we propose a self-supervised learning (SSL) framework based on past-guided future reconstruction to pretrain the motion model.
+We mask joints within the future motion sequence and compel the model to leverage the complete past sequence as guidance to reconstruct the masked elements.
+This SSL approach enables the model to acquire rich motion priors, free from the interference of a regression loss, helping it grasp the transition logic and statistical dependencies between past and future states.
+Concurrently, we also perform self-reconstruction on the past sequence to enhance the encoder’s representational power.
+To maximize the efficacy of this SSL stage, we introduce a velocity-based masking strategy.
+Unlike random masking, we prioritize masking joints that exhibit the most significant movement between adjacent frames.
+This forces the model to focus its representational capacity on the most informative and dynamic components of the motion (e.g., the upper limbs during an “eating” action), rather than on low-information, static joints.
+In the second stage, we fine-tune the pretrained model on the downstream motion prediction task.
+Here, we treat the entire future sequence as “masked” and task the model with its recovery.
+Furthermore, we explore the synergistic relationship between motion prediction and motion understanding.
+We design a novel joint fine-tuning objective where the model, given the history, is tasked with simultaneously predicting: (1) the future motion sequence (low-level prediction) and (2) a future motion description (high-level understanding).
+We find that these two tasks are mutually beneficial: the high-level, textual understanding of the future movement constrains and refines the low-level motion prediction, and vice versa.
+Our key contributions can be summarized as follows:
+  • We propose a self-supervised motion representation learning framework that unifies past and future self-reconstruction, with future reconstruction guided by historical motion cues.
+A velocity-based masking strategy is applied to emphasize dynamic joints during reconstruction.
+This paradigm enables the model to capture intrinsic spatiotemporal structure as well as the dependencies between past and future states.
+  • We develop a joint fine-tuning objective that tasks the model with predicting both the future motion sequence and future textual description.
+This coupling of low-level motion prediction and high-level semantic understanding yields mutually reinforcing improvements in coherence and realism.
   • We conduct extensive experiments and demonstrate that our method reduces the average prediction errors by 8.8% compared to state-of-the-art methods on the 3DPW, Human3.6M, and AMASS datasets.
+Moreover, with a lightweight text prediction head, our model achieves competitive performance in future motion understanding compared to LLM-based methods [wu2025mg].
 2 Related Work
+Figure 2: Overall architecture of MoReFun, which consists of a Past Motion Encoder (PME) and a Future Motion Decoder (FMD).
+The PME is built with stacked Spatial and Temporal Self-Attention layers, processing the past motion sequence.
+The FMD receives masked future motion together with the encoded past-motion features, and is composed of Spatial/Temporal Cross-Attention followed by Spatial/Temporal Self-Attention.
+Through these stages, the decoder leverages the past-motion representation to guide the reconstruction of masked future joints.
 2.1 3D Skeleton-based Human Motion Prediction
-RNN-based Methods
-Early motion prediction methods employed probability model-based approaches[38, 19].
-These methods performed poorly when dealing with complex motions.
-As RNNs can naturally process sequential data, researchers widely used RNNs and their variants for motion prediction tasks[11, 9, 14, 31].
-EDR[9] maps motion sequences to a latent space using a nonlinear encoder and models temporal features using LSTM (Long Short-Term Memory Network).
-To avoid ambiguity when predicting long-term motions, VAE-LSTM[14] introduces the encoding of control signals into a variational inference framework trained to learn human motion manifolds.
-United[31] achieves smooth prediction by modeling first-order motion derivatives.
-To incorporate spatial information modeling, Structure-RNN[18] unfolds the spatiotemporal graph in the temporal dimension and decomposes it into a set of contributing factor components fed into RNN modeling.
-SkelNet[13], at each time frame, obtains local representations of human structure through special component layer branches.
-Feed-forward Methods
-The step-by-step modeling and generation approach of RNNs often leads to the issue of error accumulation.
-In recent times, many methods have adopted a feed-forward paradigm, utilizing GNN and Transformer models, to generate prediction sequences in a non-autoregressive manner[26, 3, 10, 30, 34].
-These approaches leverage graph structures and attention computations that adapt to the non-Euclidean distribution properties of joints in space and time.
-LTD[28], HisRep[29], SGSN[22], and SPGSN[24] transform motion sequences to the frequency domain, modeling the trajectories of joints.
-DMGNN[20], MST-GNN[23], MSR-GCN[7], Sybio-GNN[21], and MMotion-Att[30] learn motion patterns at multiple levels by modeling human representations at different scales.
-DGMGNN[20], MANet[34], DD-GCN[40], and DSTD-GCN[10] model the dynamic characteristics of joints, adapting to the dynamic properties of motion.
-Transformer-based models utilize self-attention computations, which enhance the model’s capability to capture long-range dependencies between joints[1, 32, 41].
-Among them, AuxFormer[41] achieves better results by incorporating denoising and reconstruction tasks to assist the prediction task.
-In addition, lightweight models based on MLP[2, 12] have also demonstrated success.
+Early motion prediction methods relied on probabilistic models [wang2007gaussian, lehrmann2014efficient], yet their strong assumptions limited their ability to capture complex human dynamics.
+With deep learning, RNN-based approaches [ghosh2017learning, fragkiadaki2015recurrent, habibie2017recurrent, martinez2017human] became dominant by encoding historical trajectories and autoregressively generating future poses, though they often suffer from error accumulation during long-term prediction.
+To overcome this, recent works shift toward feed-forward architectures that directly predict future sequences in a non-autoregressive manner.
+Graph-based models [ma2022progressively, li2021multiscale, dang2021msr, li2021symbiotic, shi2023multi] leverage skeletal topology, frequency-domain representations [mao2019learning, mao2020history, shi2024gradient] provide compact trajectory modeling, and dynamic GNNs [li2020dynamic, fu2023learning] adapt joint relations to motion changes.
+Transformer-based models [aksan2021spatio, martinez2021pose, xu2023auxiliary] further enhance long-range dependency modeling, with AuxFormer [xu2023auxiliary] improving prediction using auxiliary denoising and reconstruction tasks.
+Meanwhile, lightweight MLP designs [bouazizi2022motionmixer, guo2023back] demonstrate that strong performance can also be achieved without heavy attention mechanisms.
+Together, these feed-forward approaches alleviate autoregressive drift and have become the mainstream paradigm for modern human motion prediction.
 2.2 Self-supervised Pretraining
 Self-supervised pretraining has brought about a revolution in the understanding and generation of language and images.
-This technique serves as one of the fundamental training methods of Large Language Models (LLMs), which is widely used in BERT[8], RoBERTa[25] and other models.
-For image processing tasks, recovering masked pixels/tokens is a common self-supervised pretraining method[15, 5].
-MAE[15] proves that even when a substantial portion of image information is masked, the network can reconstruct a complete image based on a limited number of visible features.
-The self-learning paradigm of autoencoder in MAE is also widely used in multimodal[4], video processing[36], point cloud data processing[42] and other tasks.
+This technique serves as one of the fundamental training methods of Large Language Models (LLMs), which is widely used in BERT [devlin2018bert], RoBERTa [liu2019roberta] and other models.
+For image processing tasks, recovering masked pixels/tokens is a common self-supervised pretraining method [he2022masked, chen2024context].
+MAE [he2022masked] proves that even when a substantial portion of image information is masked, the network can reconstruct a complete image based on a limited number of visible features.
+The self-learning paradigm of autoencoder in MAE is also widely used in multimodal [chen2023pimae], video processing [tong2022videomae], motion generation [shi2025genm], point cloud data processing [zhang2023learning] and other tasks.
+2.3 Motion Captioning
+Motion captioning aims to generate natural language descriptions for 3D human motion sequences [petrovich2022temos, shi2025mogic].
+Recent LLM-based approaches [li2025lamp, jiang2023motiongpt] further leverage powerful pretrained language models to produce more accurate and semantically coherent motion descriptions.
+Despite these advances, the role of future-motion semantics in enhancing motion prediction remains largely underexplored, leaving a gap between motion understanding and predicting.
 3 Method
-Figure 2: Detail of the model architecture.
-Our proposed model comprises Past Motion Encoder (PME) and Future Motion Predictor (FMP), where the former consists of $N\times$ SpatioTemporal Attention (ST-Attention), and the latter consists of $M$ ST-Attention and Past Motion Guided Attention (PMG-Attention).
-Here, the finetune stage is taken as an example to show the model structure (the pretrain stage only distinguishes the number of forward propagation of PME and incomplete input sequences).
 3.1 Problem Background
-The purpose of 3D skeleton-based human motion prediction is to generate the unseen and consecutive future motion based on past motion patterns.
-Assume a historical motion sequence consisting of $T$ frames in the Cartesian coordinate system, denoted as $\mathbf{X}^{P}=[\textbf{x}^{1},\textbf{x}^{2},...,\textbf{x}^{T}]\in\mathbb{R}^{T\times J\times K}$, where $J$ represents the number of joints and $K$ is the feature dimension (the initial value of $K$ is 3, representing the $x,y,z$ coordinates).
-Assume $\mathbf{X}^{F}=[\textbf{x}^{1},\textbf{x}^{2},...,\textbf{x}^{L}]\in\mathbb{R}^{L\times J\times K}$ is the future motion pose of $L$ frames.
-Our goal is to design and train a network $\mathcal{P}(\cdot)$ that takes $\mathbf{X}^{P}$ as input and $\mathbf{X}^{F}$ as output, i.e., $\mathbf{X}^{F}=\mathcal{P}(\mathbf{X}^{P},\pi_{\theta})$, where $\pi_{\theta}$ is the trainable parameters.
+3D skeleton-based human motion prediction aims to generate unseen and consecutive future motions from historical observations.
+Given a past motion sequence $\mathbf{X}^{P}=[\mathbf{x}_{1},\ldots,\mathbf{x}_{T}]\in\mathbb{R}^{T\times J\times K}$, where $J$ denotes the number of joints and $K$ is the feature dimension (typically $K=3$ for the $x,y,z$ coordinates), the goal is to predict the future motion sequence $\mathbf{X}^{F}=[\mathbf{x}_{1},\ldots,\mathbf{x}_{L}]\in\mathbb{R}^{L\times J\times K}$.
+We therefore design a predictive model $\mathcal{P}(\cdot)$ that maps the observed past motion to its future continuation, i.e., $\mathbf{X}^{F}=\mathcal{P}(\mathbf{X}^{P};\pi_{\theta})$ with trainable parameters $\pi_{\theta}$.
+Beyond pose prediction, we further consider the generation of a semantic description of the future behavior.
+Given the same historical sequence, the model outputs a future text sequence $\mathbf{Y}^{F}=[y_{1},\ldots,y_{D}]$ that summarizes the expected motion.
+This formulation naturally extends the task to a joint prediction problem, where the model produces both future motion and its corresponding semantic description: $(\mathbf{X}^{F},\mathbf{Y}^{F})=\mathcal{P}(\mathbf{X}^{P};\pi_{\theta})$.
+Figure 3: Overview of the finetuning stage.
+The Past Motion Encoder extracts spatiotemporal representations from the observed history motion.
+The shared past-motion feature is fed into two branches: (1) the Future Motion Decoder, which predicts the future 3D pose sequence from a zero-initialized state, and (2) the Understanding Decoder, which produces a natural-language description of the upcoming motion.
+Both branches are jointly optimized with motion and text supervision.
 3.2 Model Architecture
 As shown in Fig.
-2, our proposed model consists of a Past Motion Encoder (PME) and a Future Motion Predictor (FMP).
+2, our proposed model consists of a Past Motion Encoder (PME) and a Future Motion Decoder (FMD).
 The PME aims to extract spatiotemporal features from the past motion sequence.
-The FMP aims to generate the future motion sequence based on the initial state (which consists of zero-valued inputs in motion prediction tasks) and past motion features.
+The FMD aims to predict the future motion sequence based on the initial state (which consists of zero-valued inputs in motion prediction tasks) and past motion features.
 Joint Embedder and Predict Head
@@
 $C$ denotes the number of feature channels.
-For the output of PME and FMP, a linear layer as the Predict Head is applied to obtain the 3D motion sequence.
+For the output of PME and FMD, a linear layer as the Predict Head is applied to obtain the 3D motion sequence.
 Past Motion Encoder (PME)
@@
 Human motion sequences contain two features: the change of joint position over time and the relationship between joints in space.
-We use the SpatioTemporal Attention (ST-Attention) to simultaneously model the spatiotemporal information of the joints.
-Specifically, PME contains $N$ cascaded ST-Attention modules, where each module first models the sequence dependencies in the temporal dimension and then models the joint relationships in the spatial dimension.
-Assuming the input to the $l^{th}$ layer of ST-Transformer is $\mathbf{H}^{l}$, the computation process can be expressed as:
+We use the spatiotemporal attention to simultaneously model the spatiotemporal information of the joints.
+Specifically, PME contains $N$ cascaded spatiotemporal attention modules, where each module first models the sequence dependencies in the temporal dimension and then models the joint relationships in the spatial dimension.
+Assuming the input to the $l^{th}$ layer is $\mathbf{H}^{l}$, the computation process can be expressed as:
 $\mathbf{H}^{l+1}=\mathcal{F}_{attn}^{s}(\mathcal{F}_{attn}^{t}(\mathbf{H}^{l}))$ (2)
 where $\mathcal{F}_{attn}^{s}(\cdot)$ / $\mathcal{F}_{attn}^{t}(\cdot)$ represents the spatial/temporal multi-head self-attention operation with the bottle neck.
-Future Motion Predictor (FMP)
-The purpose of FMP is to generate future motion sequences based on the initial state and the output from PME (i.e., past motion features).
-FMP comprises two modules: the ST-Attention module for modeling the sequence dependencies within itself and the Past Movement Guided Attention module (PMG-Attention).
-PMG-Attention applies cross-attention to incorporate past motion features into the current state, which can be mathematically represented as:
+Future Motion Decoder (FMD)
+The purpose of FMD is to generate future motion sequences based on the initial state and the output from PME (i.e., past motion features).
+FMD comprises two modules: the spatiotemporal attention module for modeling the sequence dependencies within itself and the spatiotemporal cross-attention to incorporate past motion features into the current state, which can be mathematically represented as:
 $\mathbf{H}^{l+1}=\mathcal{F}_{c-attn}^{s}(\mathcal{F}_{c-attn}^{t}(\mathbf{H}^{l},\mathbf{H}^{P}),\mathbf{H}^{P})$ (3)
 where $\mathbf{H}^{P}$ is the past motion features output by PME.
-$\mathbf{H}^{l}$ is the input of $l^{th}$ PMG-Attention.
+$\mathbf{H}^{l}$ is the input of $l^{th}$ layer of FMD.
 $\mathcal{F}_{c-attn}^{s}(\cdot)$ / $\mathcal{F}_{c-attn}^{t}(\cdot)$ represents the spatial/temporal multi-head cross-attention calculation with the bottle neck.
-We provide more information about the ST-Attention and PMG-Attention in the supplementary materials (Sec.
-7).
 3.3 Past Movement Guided Pretraining
-We apply the self-reconstruction pretraining paradigm to help the model learn the spatiotemporal representation of joints in motion sequences (as shown in the left part of Fig.
-1(d)), which compromise three steps: 1) past motion reconstruction; 2) past motion encoding; 3) future motion reconstruction.
+We apply the self-reconstruction pretraining paradigm to help the model learn the spatiotemporal representation of joints in motion sequences (as shown in the Fig.
+2), which compromise two proxy tasks: 1) past motion reconstruction; 2) future motion reconstruction.
 We will introduce each step in the following.
@@
 Given the past motion sequence $\mathbf{X}^{P}\in\mathcal{R}^{T\times J\times K}$, we first calculate the motion velocity of each joint $\mathbf{V}=\mathbf{X}_{2:T}^{P}-\mathbf{X}_{1:T-1}^{P}\in\mathcal{R}^{(T-1)\times J\times K}$.
-We find the threshold $s$ in $\mathbf{V}$ such that the $r$of the values in $\mathbf{V}$ are greater than the threshold $s$.
+We find the threshold $s$ in $\mathbf{V}$ such that the rr% of the values in $\mathbf{V}$ are greater than the threshold $s$.
 The mask of the past motion sequence $\mathbf{M}^{P}$ can be mathematically expressed as:
@@
 Then, the joint embedders are applied to obtain the masked motion embeds $\mathbf{X}_{emb}^{P}$.
-Noting that we incorporate the trainable temporal and spatial position encoding into the motion embeds to learn the temporal order and spatial relationships of joints.
-Finally, we calculate the reconstructed motion sequence $\mathbf{X}_{rec}^{P}=\mathcal{F}_{PH}(\mathcal{F}_{PME}(\mathbf{X}_{emb}^{P}))$, where $\mathcal{F}_{PH}(\cdot)$ and $\mathcal{F}_{PME}(\cdot)$ are Predict Head and Position Motion Encoder.
-Past Motion Encoding
-This step aims to compute the features of complete past motion, which will be used to guide the future motion reconstruction and integrate the hidden relationship between the past sequence and the future sequence into the network.
-Given the complete input sequence $\mathbf{X}^{P}$, we first apply joint embedder to obtain the motion embeds $\mathbf{X}_{emb}^{P}$.
-Then, the past motion features can be computed by $\mathbf{H}^{P}=\mathcal{F}_{PME}(\mathbf{X}_{emb}^{P})$.
+Finally, we calculate the reconstructed motion sequence $\mathbf{X}_{rec}^{P}=\mathcal{F}_{PH}(\mathcal{F}_{PME}(\mathbf{X}_{emb}^{P}))$, where $\mathcal{F}_{PH}(\cdot)$ and $\mathcal{F}_{PME}(\cdot)$ are Predict Head and Past Motion Encoder.
+Through this past-motion self-reconstruction objective, the model learns to capture both the spatial structure within individual poses and the temporal dependencies across consecutive motions.
+This encourages the encoder to acquire rich and robust 3D human motion representations that benefit downstream tasks.
 Future Motion Reconstruction
-Given the future motion sequence $\mathbf{X}^{F}\in\mathcal{R}^{L\times J\times K}$, we first use the same method for masking and embedding the future motion sequence as for processing the past motion sequence.
-The masked future motion embeds, denoted as $\mathbf{X}_{emb}^{F}$ and the complete past motion features $\mathbf{H}^{P}$ are the input of FMP.
-In the PMG-Attention of FMP, past motion features guide the reconstruction of masked future motion through cross-attention mechanism.
-Finally, the reconstructed motion sequence can be expressed as: $\mathbf{X}_{rec}^{F}=\mathcal{F}_{PH}(\mathcal{F}_{FMP}(\mathbf{X}_{emb}^{F},\mathbf{H}^{P})))$.
-3.4 Fine-tuning for Motion Prediction
-In the finetune phase, we use the complete past motion sequence as input to the PME to get past motion features (as shown in the right part of Fig.
-1(d)).
-In FMP, we use the initial state initialized with the value of 0 (but retain the position encoding) as input, and output the complete future sequence guided by the past motion features.
-Given the past motion sequence $\mathbf{X}^{P}$, this process can be represented as:
+Given the complete past motion sequence $\mathbf{X}^{P}$, we first apply the joint embedder to obtain the motion embeddings $\mathbf{X}_{emb}^{P}$, which are then encoded into past-motion features via $\mathbf{H}^{P}=\mathcal{F}_{PME}(\mathbf{X}_{emb}^{P})$.
+For the future motion sequence $\mathbf{X}^{F}\in\mathbb{R}^{L\times J\times K}$, we adopt the same velocity-based masking strategy and embedding procedure, producing the masked embeddings $\mathbf{X}_{emb}^{F}$.
+The masked future-motion embeddings $\mathbf{X}_{emb}^{F}$ and the past-motion features $\mathbf{H}^{P}$ are fed into the FMD module, where the PMG-Attention uses cross-attention to guide the reconstruction of masked future motion.
+The final reconstructed future motion is given by $\mathbf{X}_{rec}^{F}=\mathcal{F}_{PH}(\mathcal{F}_{FMD}(\mathbf{X}_{emb}^{F},\mathbf{H}^{P}))$.
+In this reconstruction task, the model completes masked future motions based on past observations.
+This encourages it to explicitly learn the fundamental dependency of future motion on past inputs, as well as the underlying priors of human movement patterns and motion trajectories, facilitating downstream tasks.
+3.4 Fine-tuning for Prediction and Understanding
+In the finetuning phase, we use the complete past motion sequence as input to the PME to obtain the past motion features (as shown in the Fig.3).
+In the FMD, we initialize the future query tokens with zeros (while preserving positional encodings) and decode the full future motion sequence guided by the past features.
+Given the past motion sequence $\mathbf{X}^{P}$, the prediction process is formulated as:
 $\mathbf{H}^{P}$ $=\mathcal{F}_{JE}(\mathcal{F}_{PME}(\mathbf{X}^{P}))$ (5)
-$\mathbf{H}^{F}$ $=\mathcal{F}_{JE}(\mathcal{F}_{FMP}(\mathbf{H}_{0}^{F},\mathbf{H}^{P}))$ (6)
-$\mathbf{X}^{F}$ $=\mathcal{F}_{PH}(\mathbf{H}^{F})$ (7)
-where $\mathbf{H}_{0}^{F}$ is the initial state initialized with the value of 0. $\mathcal{F}_{JE}$ represents the joint embedder.
+$\mathbf{H}^{F}$ $=\mathcal{F}_{JE}(\mathcal{F}_{FMD}(\mathbf{H}_{0}^{F},\mathbf{H}^{P}))$ (6)
+$\mathbf{X}^{F}$ $=\mathcal{F}_{PH}(\mathbf{H}^{F}),$ (7)
+where $\mathbf{H}_{0}^{F}$ is the zero-initialized future state and $\mathcal{F}_{JE}$ denotes the joint embedder.
+Beyond motion prediction, our framework also supports future motion understanding.
+Given the past-context representation $\mathbf{H}^{P}$, we apply a Future Semantic Decoder (FSD) $\mathcal{F}_{FSD}$ to generate a future semantic description:
+$\mathbf{Y}^{F}=\mathcal{F}_{FSD}(\mathbf{H}^{F},\mathbf{H}^{P}),$ (8)
+where $\mathbf{Y}^{F}=[y_{1},\ldots,y_{M}]$ is the predicted future text sequence.
+This enables the model to jointly perform future motion prediction and future motion understanding within a unified fine-tuning pipeline.
 3.5 Loss Function
-Loss Function of Pretrain Stage
-Pretrain stage performs past sequence reconstruction and future sequence reconstruction.
+Loss Function of Pretraining Stage
+Pretraining stage performs past sequence reconstruction and future sequence reconstruction.
 Given the past and future reconstructed results $\hat{\mathbf{X}}_{rec}^{P}$ and $\hat{\mathbf{X}}_{rec}^{F}$, and the corresponding ground-truth (i.e.
 complete sequence) $\mathbf{X}_{rec}^{P}$ and $\mathbf{X}_{rec}^{F}$, the loss function can be represented as follow:
-$\begin{split}\mathcal{L}_{pretrain}&=\frac{1}{TJ}\sum_{t=1}^{T}\sum_{j=1}^{J}||\mathbf{X}_{rec}^{P,(t,j)}-\hat{\mathbf{X}}_{rec}^{P,(t,j)}||_{F}^{2}\\ &+\frac{\alpha}{LJ}\sum_{l=1}^{L}\sum_{j=1}^{J}||\mathbf{X}_{rec}^{F,(l,j)}-\hat{\mathbf{X}}_{rec}^{F,(l,j)}||_{F}^{2}\end{split}$ (8)
+[…]
 where $\alpha$ is the hyperparameter to balance the loss weight.
 Loss Function of Finetune Stage
-For finetune stage, given the predicted sequence $\hat{\mathbf{X}}$ and the ground-truth $\mathbf{X}$, the loss function can be defined as:
-$\mathcal{L}_{finetune}=\frac{1}{LJ}\sum_{l=1}^{L}\sum_{j=1}^{J}||\mathbf{X}_{j}^{l}-\hat{\mathbf{X}}_{j}^{l}||_{F}^{2}$ (9)
+In the finetuning stage, the overall objective integrates motion reconstruction and semantic prediction into a single loss:
+$\mathcal{L}_{finetune}$ $=\underbrace{\frac{1}{LJ}\sum_{l=1}^{L}\sum_{j=1}^{J}\left\|\mathbf{X}_{j}^{l}-\hat{\mathbf{X}}_{j}^{l}\right\|_{F}^{2}}_{\mathcal{L}_{motion}}$ (10)
+$\quad+\lambda\underbrace{\left(-\sum_{d=1}^{D}\log p(\mathbf{Y}^{F}_{d}\mid\hat{\mathbf{Y}}^{F})\right)}_{\mathcal{L}_{text}}.$
+Jointly optimizing the two losses encourages mutually reinforcing representations, leading to more accurate future motion prediction and more consistent semantic generation.
 Table 1: Comparisons of MPJPEs between our proposed method with other state-of-the-art methods for short-term prediction on Human3.6.
@@
 Milliseconds 80 160 320 400 80 160 320 400 80 160 320 400 80 160 320 400
-DMGNN[20] 17.3 30.7 54.6 65.2 11.0 21.4 36.2 43.9 9.0 17.6 32.1 40.3 17.3 34.8 61.0 69.8
-MSR[7] 12.2 22.7 38.6 45.2 8.4 17.1 33.0 40.4 8.0 16.3 31.3 38.2 12.0 26.8 57.1 69.7
-Mixer[2] 12.0 23.0 42.4 51.3 7.8 16.5 34.3 43.0 7.4 15.2 30.3 37.6 10.7 25.0 56.9 71.1
-PGBIG[26] 10.2 19.8 34.5 40.3 7.0 15.1 30.6 38.1 6.6 14.1 28.2 34.7 10.0 23.8 53.6 66.7
-SPGSN[24] 10.1 19.4 34.8 41.5 7.1 14.8 30.5 37.9 6.7 13.8 28.0 34.6 10.4 23.8 53.6 67.1
-siMLPe[12] 11.9 19.3 31.9 36.9 10.2 16.5 30.7 37.3 10.4 16.4 29.7 35.9 12.7 25.3 55.1 68.1
-DSTD-GC[10] 11.1 22.4 38.8 45.2 7.0 15.5 31.7 39.2 6.6 14.8 29.8 36.7 10.0 24.4 54.5 67.4
-AuxFormer[41] 8.9 16.9 30.1 36.1 6.4 14.0 28.8 35.9 5.7 11.4 22.1 27.9 8.6 18.8 38.8 49.2
-CIST-GCN[33] 11.8 23.4 40.5 46.5 6.7 14.8 29.8 36.8 7.3 15.6 31.0 38.0 10.2 23.7 52.3 65.3
+DSTD-GC [fu2023learning] 11.1 22.4 38.8 45.2 7.0 15.5 31.7 39.2 6.6 14.8 29.8 36.7 10.0 24.4 54.5 67.4
+AuxFormer [xu2023auxiliary] 8.9 16.9 30.1 36.1 6.4 14.0 28.8 35.9 5.7 11.4 22.1 27.9 8.6 18.8 38.8 49.2
+CIST-GCN [medina2024context] 11.8 23.4 40.5 46.5 6.7 14.8 29.8 36.8 7.3 15.6 31.0 38.0 10.2 23.7 52.3 65.3
 Ours 8.6 16.5 31.2 37.7 6.1 13.7 29.3 36.8 5.3 10.3 19.7 25.2 7.7 16.3 31.7 40.9
@@
 Milliseconds 80 160 320 400 80 160 320 400 80 160 320 400 80 160 320 400
-DMGNN[20] 13.1 24.6 64.7 81.9 23.3 50.3 107.3 132.1 12.5 25.8 48.1 58.3 15.3 29.3 71.5 96.7
-MSR[7] 8.6 19.7 43.3 53.8 16.5 37.0 77.3 93.4 10.1 20.7 41.5 51.3 12.8 29.4 67.0 85.0
-Mixer[2] 7.5 18.4 43.9 55.3 16.4 37.1 80.2 98.2 9.1 19.5 42.4 53.6 11.5 27.0 64.9 83.5
-PGBIG[26] 7.2 17.6 40.9 51.5 15.2 34.1 71.6 87.1 8.3 18.3 38.7 48.4 10.7 25.7 60.0 76.6
-SPGSN[24] 7.4 17.2 40.0 50.3 14.6 32.6 70.6 86.4 8.7 18.3 38.7 48.5 10.7 25.3 59.9 76.5
-siMLPe[12] 11.0 19.4 41.6 52.1 16.2 33.7 70.9 86.3 11.9 20.3 39.7 48.9 13.9 27.3 62.0 78.6
-DSTD-GC[10] 6.9 17.4 40.1 51.7 14.3 33.5 72.2 87.3 8.5 19.2 40.3 49.9 10.1 25.4 60.6 77.3
-AuxFormer[41] 6.8 17.0 40.3 51.6 13.5 31.3 69.2 85.4 7.9 17.3 37.4 47.2 8.8 19.1 39.2 51.0
-CIST-GCN[33] 7.3 18.1 43.6 55.3 13.7 31.0 65.7 79.9 8.6 18.5 39.3 49.6 9.6 23.7 57.7 75.0
+DSTD-GC [fu2023learning] 6.9 17.4 40.1 51.7 14.3 33.5 72.2 87.3 8.5 19.2 40.3 49.9 10.1 25.4 60.6 77.3
+AuxFormer [xu2023auxiliary] 6.8 17.0 40.3 51.6 13.5 31.3 69.2 85.4 7.9 17.3 37.4 47.2 8.8 19.1 39.2 51.0
+CIST-GCN [medina2024context] 7.3 18.1 43.6 55.3 13.7 31.0 65.7 79.9 8.6 18.5 39.3 49.6 9.6 23.7 57.7 75.0
 Ours 6.3 15.8 39.8 51.4 10.8 26.4 62.7 79.8 7.5 17.1 37.9 48.0 8.2 16.9 32.8 43.5
@@
 The best results are highlighted in bold.
-Motion Average Param(M)
-Milliseconds 80 160 320 400 560 1000 -
-DMGNN[20] 17.0 33.6 65.9 79.7 103.0 137.2 4.82
-MSR[7] 12.1 25.6 51.6 62.9 81.1 114.2 6.30
-Mixer[2] 11.1 24.0 51.5 64.0 83.5 116.5 0.03
-PGBIG[26] 10.3 22.7 47.4 58.5 76.9 110.3 1.80
-SPGSN[24] 10.4 22.3 47.1 58.3 77.1 108.8 5.66
-siMLPe[12] 13.4 23.7 47.4 58.0 77.9 111.7 0.14
-DSTD-GC[10] 10.4 23.3 48.8 59.8 77.8 111.0 0.18
-AuxFormer[41] 9.5 20.6 43.4 54.1 75.3 107.0 1.00
-CIST-GCN[33] 10.5 23.2 47.9 59.0 77.2 110.3 0.35
-GCNext [39] 9.3 21.5 45.5 56.4 74.7 108.7 -
-Ours 8.8 19.3 41.1 51.8 70.0 101.5 1.66
+Motion Average
+Milliseconds 80 160 320 400 560 1000
+DMGNN [li2020dynamic] 17.0 33.6 65.9 79.7 103.0 137.2
+MSR [dang2021msr] 12.1 25.6 51.6 62.9 81.1 114.2
+Mixer [bouazizi2022motionmixer] 11.1 24.0 51.5 64.0 83.5 116.5
+PGBIG [ma2022progressively] 10.3 22.7 47.4 58.5 76.9 110.3
+SPGSN [li2022skeleton] 10.4 22.3 47.1 58.3 77.1 108.8
+siMLPe [guo2023back] 13.4 23.7 47.4 58.0 77.9 111.7
+DSTD-GC [fu2023learning] 10.4 23.3 48.8 59.8 77.8 111.0
+AuxFormer [xu2023auxiliary] 9.5 20.6 43.4 54.1 75.3 107.0
+CIST-GCN [medina2024context] 10.5 23.2 47.9 59.0 77.2 110.3
+GCNext [wang2024gcnext] 9.3 21.5 45.5 56.4 74.7 108.7
+ALIEN [wei2025alien] 9.7 21.8 47.0 58.3 - -
+Ours 8.8 19.3 41.1 51.8 70.0 101.5
 4 Experiments
@@
 Human3.6M
-Human3.6M[17] is a large-scale, accurately annotated 3D human pose dataset that captures the motions of 11 professional actors in daily life scenarios using motion capture technology, totaling over 3.6 million pose instances.
-Following previous works[26, 24], only 22 joints are used, focusing on those that exhibit significant movement, and all sequences are down-sampled to 25 fps.
+Human3.6M [ionescu2013human3] is a large-scale, accurately annotated 3D human pose dataset that captures the motions of 11 professional actors in daily life scenarios using motion capture technology, totaling over 3.6 million pose instances.
+Following previous works [ma2022progressively, li2022skeleton], only 22 joints are used, focusing on those that exhibit significant movement, and all sequences are down-sampled to 25 fps.
 The 3D Poses in the Wild (3DPW)
-3DPW[37] stands out as a pioneering dataset in motion related tasks, primarily due to its focus on capturing human motion in natural settings outside of the typical laboratory environment.
-This dataset provides a rich collection of video sequences captured at a frame rate of 30 frames per second, which includes diverse activities such as ’Walking’, ’Running’, and interacting with various objects.
+3DPW [von2018recovering] stands out as a pioneering dataset in motion related tasks, primarily due to its focus on capturing human motion in natural settings outside of the typical laboratory environment.
+This dataset provides a rich collection of video sequences captured at a frame rate of 30 frames per second, which includes diverse activities such as ‘Walking’, ‘Running’, and interacting with various objects.
 The Archive of Motion Capture as Surface Shapes (AMASS)
-AMASS[27] consolidates 15 existing motion capture datasets into a single, standardized framework.
+AMASS [mahmood2019amass] consolidates 15 existing motion capture datasets into a single, standardized framework.
 By converting these diverse datasets into a common representation, AMASS provides over 40 hours of meticulously annotated motion data.
-Following the previous works[34, 40], we discarded the detailed joints of the hand and sampled the sequence to 24fps.
-Figure 3: Visualization results of short-term prediction (a) and reconstruction (b).
-The blue lines represent predicted/reconstructed motions, and the red lines represent ground-truths.
-We map the results to 2D space for display.
+Following the previous works [shi2023multi, wang2024dynamic], we discarded the detailed joints of the hand and sampled the sequence to 24fps.
+FineMotion
+To incorporate fine-grained action semantics, we adopt FineMotion [^1] https://github.com/CVI-SZU/FineMotion, an open-source extension of HumanML3D that re-labels all motions with detailed body-part movement descriptions.
+FineMotion segments each sequence into 0.5-second snippets and provides both human-annotated and automatically generated descriptions.
+Based on this resource, we construct our prediction dataset using a 2-second sliding window (0.5s past motion and 1.5s future motion).
+For each window, we concatenate the snippet-level descriptions corresponding to the future segment to form its semantic target.
+This process yields 12,628 human-annotated samples and 248,126 automatically annotated samples, which we split into train/val/test sets following an 80% / 5% / 15% ratio.
+Figure 4: Visualization results of motion prediction. The red lines represent predicted motions, and the blue lines represent ground-truths.
+Figure 5: Visualization results of motion reconstruction on H36M. The blue lines represent reconstructed motions, and the red lines represent ground-truths.
+Figure 6: Visualization of motion prediction and understanding.
 4.2 Experimental Settings
-For the PME and the FMP, the number of modules (denoted as $N,M$) are both set to 3.
-The number of feature channels remains consistent at 128 throughout the network architecture.
+For the PME, FMD and FSU, the number of modules are all set to 3.
+The number of feature channels remains consistent at 128 throughout the PME and FMD, and remains at 256 throughout the FSU.
 In the attention calculation, we employed 8 heads, with each head comprising 32 feature channels.
 During the pretraining phase, the loss weight (denoted as $\alpha$) is set to 1, indicating an equal contribution of loss from both past and future sequence reconstruction.
-We provide more implementation details in the supplementary materials (Sec.
-8).
+During the finetune phase, the loss weight (denoted as $\lambda$) is set to 0.1.
+The parameters in FSU are updated every 5 epoch.
 4.3 Comparison to State-of-the-art Methods
-We compare our method with the recent state-of-the-art methods, including DMGNN (CVPR’20)[20, 30, 1, 7, 2, 26, 24, 12, 10, 41, 40, 33, 39].
+We compare our method with the recent state-of-the-art methods [li2020dynamic, mao2021multi, aksan2021spatio, dang2021msr, bouazizi2022motionmixer, ma2022progressively, li2022skeleton, guo2023back, fu2023learning, xu2023auxiliary, wang2024dynamic, medina2024context, wang2024gcnext, wei2025alien].
 To ensure fair comparisons, we re-evaluated some methods on the entire test set to align the testing standard using their official pretrained models, or retrain the model (Only for the cases where the official pre-training model is not provided) using their default hyperparameters on the Human3.6M.
 Comparisons on Human3.6M
-We use MPJPE (Mean Per Joint Position Error) as the evaluation metric, which is widely used in motion prediction tasks[26, 24, 41].
+We use MPJPE (Mean Per Joint Position Error) as the evaluation metric, which is widely used in motion prediction tasks [ma2022progressively, li2022skeleton, xu2023auxiliary].
 Tab.
@@
 Detailed actions are listed.
-As can be seen, our method outperforms the others on most actions, especially the ”Smoking”, ”Greeting”, and ”Posing”.
+As can be seen, our method outperforms the others on most actions, especially the “Smoking”, “Greeting”, and “Posing”.
 Tab.
@@
 Our approach effectively reduces prediction errors by 5.4%/10.2%/9.7%/8.2%/6.3%/6.6% at 80ms/160ms/320ms/400ms/560ms/1000ms.
-Table 3: Comparisons of MPJPEs between our proposed method with other state-of-the-art methods for both short-term and long-term prediction on AMASS. The best results are highlighted in bold.
+Table 3: Comparisons of MPJPEs between our proposed method with other SOTA methods for both short-term and long-term prediction on AMASS. The best results are highlighted in bold.
 Milliseconds 80 160 320 400 560 720 880 1000
-LTD-10-10[28] 10.3 19.3 36.6 44.6 61.5 75.9 86.2 91.2
-LTD-10-25[28] 11.0 20.7 37.8 45.3 57.2 65.7 71.3 75.2
-HisRep[29] 11.3 20.7 35.7 42.0 51.7 58.6 63.4 67.2
-Mao W, et al.[30] 11.0 20.3 35.0 41.2 50.7 57.4 61.9 65.8
-siMLPe[12] 10.8 19.6 34.3 40.5 50.5 57.3 62.4 65.7
-DD-GCN[40] 10.4 19.1 33.6 39.8 49.3 56.5 61.3 64.6
+LTD-10-10 [mao2019learning] 10.3 19.3 36.6 44.6 61.5 75.9 86.2 91.2
+LTD-10-25 [mao2019learning] 11.0 20.7 37.8 45.3 57.2 65.7 71.3 75.2
+HisRep [mao2020history] 11.3 20.7 35.7 42.0 51.7 58.6 63.4 67.2
+Mao W, et al. [mao2021multi] 11.0 20.3 35.0 41.2 50.7 57.4 61.9 65.8
+siMLPe [guo2023back] 10.8 19.6 34.3 40.5 50.5 57.3 62.4 65.7
+DD-GCN [wang2024dynamic] 10.4 19.1 33.6 39.8 49.3 56.5 61.3 64.6
 Ours 8.6 15.3 27.7 33.2 42.0 48.3 53.0 56.0
@@
 3 shows the comparisons on AMASS dataset.
-Our method achieves a reduction in prediction errors of 17.3%/19.9%/17.6%/16.6%/14.8%/14.5%/13.5%/13.3% at 80ms/160ms/320ms/400ms/560ms/720ms/880ms/1000ms.
+Our method achieves a reduction in prediction errors of 15.9%.
 We found that our method exhibits the most significant improvement on this dataset.
@@
 Milliseconds 100 200 400 600 800 1000
-DMGNN[20] 17.80 37.11 70.38 94.12 109.67 123.93
-HisRep[29] 15.88 35.14 66.82 93.55 107.63 114.75
-MSR-GCN[7] 15.70 33.48 65.02 93.81 108.15 116.31
-STSGCN[35] 18.32 37.79 67.51 92.75 106.65 112.22
-PGBIG[26] 17.66 35.32 67.83 89.60 102.59 109.41
-SPGSN[24] 15.39 32.91 64.54 91.62 103.98 111.05
-AuxFormer[41] 14.21 30.04 58.50 89.45 100.78 107.45
+DMGNN [li2020dynamic] 17.80 37.11 70.38 94.12 109.67 123.93
+HisRep [mao2020history] 15.88 35.14 66.82 93.55 107.63 114.75
+MSR-GCN [dang2021msr] 15.70 33.48 65.02 93.81 108.15 116.31
+PGBIG [ma2022progressively] 17.66 35.32 67.83 89.60 102.59 109.41
+SPGSN [li2022skeleton] 15.39 32.91 64.54 91.62 103.98 111.05
+AuxFormer [xu2023auxiliary] 14.21 30.04 58.50 89.45 100.78 107.45
+ALIEN [wei2025alien] 14.83 32.59 62.06 90.25 103.62 -
 Ours 13.85 29.11 57.06 87.35 97.93 103.37
@@
 4, our approach performed best across all frames, showcasing its robustness in handling complex interactive scenarios.
-Table 5: Comparison of Motion Classification Using Network-Output Features
-w/o Network AuxFormer[41] Ours
-Accuracy 25.6% 80.8% 95.0%
+Comparisons on Joint Prediction and Understading
+As shown in Tab.
+6, our method consistently achieves lower MPJPEs on the processed FineMotion dataset across all prediction horizons compared to prior approaches.
+Notably, introducing the understanding branch further improves motion prediction quality, indicating that capturing semantic cues of future movements can, in turn, benefit joint-level forecasting.
+As shown in Tab.
+7, the IDs “A” and “B” refer to models trained on auto-labeled and human-labeled text annotations, respectively.
+Our lightweight semantic prediction head achieves competitive or superior performance across several key metrics when compared with MGMotionLLM [wu2025mg], despite the latter being a large-scale language model designed for motion–text generation.
+Moreover, when the motion prediction loss is removed, the performance on future text generation also degrades, indicating that accurate motion forecasting provides essential guidance for understanding future semantic.
+Table 5: Comparison of classification accuracy and clustering quality on Human3.6M using network-output features.
+CLS accuracy evaluates discriminativeness, while inter- and intra-class distances measure feature separability.
+Task Metrics w/o Network AuxFormer [xu2023auxiliary] Ours
+[…]
+[…]
+[…]
 Representation Capability
-To validate the representation learning capability of our proposed method, we first extract motion features from the outputs of the trained model (outputs from the penultimate layer).
-Subsequently, we train a simple linear classifier to classify these features.
-As shown in Tab.
-5, our method significantly improves classification accuracy compared to methods that do not involve feature extraction via neural networks and AuxFormer[41].
+We evaluate the quality of the learned motion representation by extracting penultimate-layer features and training a lightweight linear classifier on them.
+As shown in Table 5, our method achieves substantially higher classification accuracy than both w/o Network and AuxFormer [xu2023auxiliary], demonstrating stronger feature discriminativeness.
+In addition, our features present larger inter-class distances and smaller intra-class distances, indicating better clustering behavior and clearer semantic separability.
+For the clustering metrics, we randomly sample 100 instances from each of the 15 action categories in Human3.6M to compute the intra- and inter-class distances.
 4.4 Visualization of Prediction and Reconstruction
 Fig.
-3 (a) presents the visualization of prediction results.
-We conducted comparisons with PGBIG[26] and Mixer[2] on challenging samples.
-In the ’Posing’ and ’Discussion’ actions, our model successfully captures significant upward/downward movements of the arms.
+4 presents the visualization of prediction results.
+We conducted comparisons with GCNext [wang2024gcnext] on challenging samples.
+Our model successfully captures significant movements of the arms and legs.
 This can be attributed to the utilization of past motion guidance, which enables the model to comprehend the influence of previous motion patterns on future sequences.
-We also provide reconstruction results with a mask rate of 75% in Fig.
-3 (b).
-Similar to the property of mask-based pretraining[15], even with a high mask rate of 75% for joints, the model can still accurately reconstruct the exact location of them.
+Fig.
+6 showcases the joint results of future motion prediction and future motion understanding, where our method consistently outputs motion sequences and semantic descriptions that are highly aligned.
+Furthermore, We provide reconstruction results with a mask rate of 75% in Fig.
+5.
+Similar to the property of mask-based pretraining [he2022masked], even with a high mask rate of 75% for joints, the model can still accurately reconstruct the location of them.
+Table 6: Comparisons of MPJPEs between our proposed method with other state-of-the-art methods on the processed FineMotion dataset. The best results are highlighted in bold.
+Average MPJPE
+Milliseconds 300 600 900 1200 1500
+AuxFormer [xu2023auxiliary] 36.07 62.31 76.95 86.11 94.22
+GCNext [wang2024gcnext] 44.05 73.19 88.14 97.03 103.9
+Ours w/o Und. 34.96 57.39 69.08 76.58 84.95
+Ours 34.69 56.71 67.99 75.33 83.65
+Table 7: Comparisons of future motion understanding between our proposed method with MGMotionLLM [wu2025mg] on the processed FineMotion dataset. The best results are highlighted in bold.
+ID Method BLEU@4 BLEU@1 ROUGE CIDEr
+A MGMotionLLM [wu2025mg] 25.3 15.8 38.0 12.5
+Ours w/o Prediction 15.6 7.9 31.2 36.2
+Ours 17.1 9.4 33.5 47.6
+B MGMotionLLM [wu2025mg] 13.1 7.5 35.8 3.3
+Ours w/o Prediction 15.0 7.1 31.7 25.8
+Ours 18.2 7.8 32.5 24.2
 5 Ablation Study
-In this section, we explore the impact of components of our network, including the mask strategy, pretraining paradigm.
+We explore the impact of components of our network, including the mask strategy, pretraining paradigm.
 We provide studies on model architecture in the supplementary materials.
@@
 Tab.
-6 also shows the comparisons of different mask rates (ranging from 0% to 75%), represented by B, C, D, F.
+8 also shows the comparisons of different mask rates (ranging from 25% to 75%), represented by B, C, F.
 The high mask rate facilitates the model to learn the space-time representation during the self-reconstruction process.
@@
 Compared with the self-reconstruction, denoising pretraining method lacks the exploration of the spatiotemporal relationship between joints.
-Table 6: Impact of mask strategy and pretraining paradigm. $r$ is the mask rate. ’denoise’ means we conduct self-supervised pretraining by denoising.
+The result without pretraining is denoted as D, which clearly demonstrates the crucial role of our pretraining strategy in improving prediction performance.
+Table 8: Impact of mask strategy and pretraining paradigm. $r$ is the mask rate. “denoise” means we conduct self-supervised pretraining by denoising.
 ID 80ms 160ms 320ms 400ms
 A random mask 9.0 19.7 42.4 52.8
-B $r=0\%$ 9.6 21.9 46.2 56.8
-C $r=25\%$ 9.3 20.2 42.9 53.7
-D $r=50\%$ 9.1 20.0 42.3 52.6
+B $r=25\%$ 9.3 20.2 42.9 53.7
+C $r=50\%$ 9.1 20.0 42.3 52.6
+D w/o pretraining 9.6 21.9 46.2 56.8
 E denoise 9.2 20.4 42.9 53.8
@@
 6 Conclusion
-In this paper, we propose a novel motion prediction paradigm, which consists two stages: 1) Past motion guided future motion self-reconstruction for learning the spatio-temporal representation of human joints (i.e.
-pretrain stage).
-2) Train the network for the motion prediction task (i.e.
-finetune stage).
-Despite the simplicity of the model structure, our approach achieves the best results on three representative datasets and is highly extensible.
+In this work, we addressed the core limitations of end-to-end regression in motion prediction by introducing a two-stage framework that separates representation learning from prediction.
+Through past-guided future reconstruction and velocity-based masking, our self-supervised pretraining effectively captures intrinsic spatiotemporal dependencies.
+The joint fine-tuning of motion prediction and semantic future understanding further enhances coherence and realism.
+[…]
 Supplementary Material
-Table 7: Comparisons of MPJPEs between our proposed method with other state-of-the-art methods for short-term prediction on Human3.6.
-We show the results across all actions such as walking and eating.
-The best results are highlighted in bold, and the second best results is underlined
-Motion Walking Eating Smoking Discussion Direction
-Milliseconds 80 160 320 400 80 160 320 400 80 160 320 400 80 160 320 400 80 160 320 400
-DMGNN[20] 17.3 30.7 54.6 65.2 11.0 21.4 36.2 43.9 9.0 17.6 32.1 40.3 17.3 34.8 61.0 69.8 13.1 24.6 64.7 81.9
-MSR[7] 12.2 22.7 38.6 45.2 8.4 17.1 33.0 40.4 8.0 16.3 31.3 38.2 12.0 26.8 57.1 69.7 8.6 19.7 43.3 53.8
-Mixer[2] 12.0 23.0 42.4 51.3 7.8 16.5 34.3 43.0 7.4 15.2 30.3 37.6 10.7 25.0 56.9 71.1 7.5 18.4 43.9 55.3
-PGBIG[26] 10.2 19.8 34.5 40.3 7.0 15.1 30.6 38.1 6.6 14.1 28.2 34.7 10.0 23.8 53.6 66.7 7.2 17.6 40.9 51.5
-SPGSN[24] 10.1 19.4 34.8 41.5 7.1 14.8 30.5 37.9 6.7 13.8 28.0 34.6 10.4 23.8 53.6 67.1 7.4 17.2 40.0 50.3
-siMLPe[12] 11.9 19.3 31.9 36.9 10.2 16.5 30.7 37.3 10.4 16.4 29.7 35.9 12.7 25.3 55.1 68.1 11.0 19.4 41.6 52.1
-DSTD-GC[10] 11.1 22.4 38.8 45.2 7.0 15.5 31.7 39.2 6.6 14.8 29.8 36.7 10.0 24.4 54.5 67.4 6.9 17.4 40.1 51.7
-AuxFormer[41] 8.9 16.9 30.1 36.1 6.4 14.0 28.8 35.9 5.7 11.4 22.1 27.9 8.6 18.8 38.8 49.2 6.8 17.0 40.3 51.6
-CIST-GCN[33] 11.8 23.4 40.5 46.5 6.7 14.8 29.8 36.8 7.3 15.6 31.0 38.0 10.2 23.7 52.3 65.3 7.3 18.1 43.6 55.3
-Ours 8.6 16.5 31.2 37.7 6.1 13.7 29.3 36.8 5.3 10.3 19.7 25.2 7.7 16.3 31.7 40.9 6.3 15.8 39.8 51.4
-Motion Greeting Phoning Posing Purchases Sitting
-Milliseconds 80 160 320 400 80 160 320 400 80 160 320 400 80 160 320 400 80 160 320 400
-DMGNN[20] 23.3 50.3 107.3 132.1 12.5 25.8 48.1 58.3 15.3 29.3 71.5 96.7 21.4 38.7 75.7 92.7 11.9 25.1 44.6 50.2
-MSR[7] 16.5 37.0 77.3 93.4 10.1 20.7 41.5 51.3 12.8 29.4 67.0 85.0 14.8 32.4 66.1 79.6 10.5 22.0 46.3 57.8
-Mixer[2] 16.4 37.1 80.2 98.2 9.1 19.5 42.4 53.6 11.5 27.0 64.9 83.5 13.4 30.7 65.6 80.0 9.8 20.9 45.6 57.3
-PGBIG[26] 15.2 34.1 71.6 87.1 8.3 18.3 38.7 48.4 10.7 25.7 60.0 76.6 12.5 28.7 60.1 73.3 8.8 19.2 42.4 53.8
-SPGSN[24] 14.6 32.6 70.6 86.4 8.7 18.3 38.7 48.5 10.7 25.3 59.9 76.5 12.7 28.6 61.0 74.4 9.3 19.4 42.2 53.6
-siMLPe[12] 16.2 33.7 70.9 86.3 11.9 20.3 39.7 48.9 13.9 27.3 62.0 78.6 15.5 29.3 59.3 72.0 13.2 21.9 44.5 55.7
-DSTD-GC[10] 14.3 33.5 72.2 87.3 8.5 19.2 40.3 49.9 10.1 25.4 60.6 77.3 12.7 29.7 62.3 75.8 8.8 19.3 42.9 54.3
-AuxFormer[41] 13.5 31.3 69.2 85.4 7.9 17.3 37.4 47.2 8.8 19.1 39.2 51.0 11.9 28.0 61.8 76.3 8.7 19.0 42.1 53.3
-CIST-GCN[33] 13.7 31.0 65.7 79.9 8.6 18.5 39.3 49.6 9.6 23.7 57.7 75.0 13.3 30.2 63.0 77.3 8.9 19.4 42.3 53.6
-Ours 10.8 26.4 62.7 79.8 7.5 17.1 37.9 48.0 8.2 16.9 32.8 43.5 11.4 27.6 62.6 77.6 8.5 19.2 44.3 56.7
-Motion Sitting Down Taking Photo Waiting Walking Dog Walking Together
-Milliseconds 80 160 320 400 80 160 320 400 80 160 320 400 80 160 320 400 80 160 320 400
-DMGNN[20] 15.0 32.9 77.1 90.3 13.6 29.0 46.0 58.8 12.2 24.2 59.6 77.5 47.1 93.3 160.1 171.2 14.3 26.7 50.1 63.2
-MSR[7] 16.1 31.6 62.5 76.8 9.9 21.0 44.6 56.3 10.7 23.1 48.3 59.2 20.7 42.9 80.4 93.3 10.6 20.9 37.4 43.9
-Mixer[2] 15.3 30.5 62.7 78.0 9.0 19.9 44.7 56.9 9.9 22.0 49.7 62.5 21.1 43.9 84.7 101.2 9.9 20.4 39.3 47.3
-PGBIG[26] 13.9 27.9 57.4 71.5 8.4 18.9 42.0 53.3 8.9 20.1 43.6 54.3 18.8 39.3 73.7 86.4 8.7 18.6 34.4 41.0
-SPGSN[24] 14.2 27.7 56.7 70.7 8.8 18.9 41.5 52.7 9.2 19.8 43.1 54.1 - - - - 8.9 18.2 33.8 40.9
-siMLPe[12] 18.0 29.8 57.7 71.2 13.1 21.3 42.9 53.6 11.9 21.5 44.4 55.0 21.8 40.6 74.6 87.0 11.9 19.8 34.2 39.8
-DSTD-GC[10] 14.1 28.0 57.3 71.2 8.4 18.8 42.0 53.5 8.7 20.2 44.3 55.3 19.6 41.8 77.6 90.2 9.1 19.8 36.3 42.7
-AuxFormer[41] 13.5 27.6 57.7 72.2 8.2 18.4 41.5 53.0 8.2 18.5 41.2 52.2 17.1 36.5 70.4 83.0 7.8 15.9 30.2 37.0
-CIST-GCN[33] 14.1 29.8 57.3 69.8 8.2 18.4 40.6 51.8 8.6 19.4 43.5 54.8 20.0 41.4 73.7 85.1 9.6 20.3 38.2 45.6
-Ours 10.1 17.2 30.2 39.2 7.6 17.8 42.0 54.1 7.8 18.0 41.6 53.2 19.1 41.1 80.6 95.2 7.3 15.3 30.8 37.9
-In this supplementary material, we provide more information that could not be included in the main manuscript because of space limit, including explanation of ST-Attention and PMG-Attention (in Sec.
-7), implementation details (in Sec.
-8), complete results for testing on Human3.6M dataset (in Sec.
-9), ablation study on model architecture (in Sec.
-10), and visualization of feature map (in Sec.
-11).
-7 Detailed Explanation of ST-Attention and PMG-Attention
-In our model architecture for human motion prediction, both the Past Motion Encoder (PME) and the Future Motion Predictor (FMP) modules rely on attention mechanisms: SpatioTemporal Attention (ST-Attention) and Past Movement Guided Attention (PMG-Attention).
-These mechanisms are pivotal for capturing and integrating spatiotemporal dependencies in motion sequences.
-ST-Attention
-ST-Attention combines temporal and spatial attention mechanisms to model joint relationships over time effectively. This attention mechanism is utilized within PME to encode past motion features.
-The temporal attention mechanism operates on the sequence dependencies over time in the joint positions:
-$\mathbf{Q}^{t},\mathbf{K}^{t},\mathbf{V}^{t}=\mathcal{F}^{t}_{QKV}\left(\mathbf{H}^{l}\right)\$ (10)
-$\mathbf{Z}^{t}=\operatorname{softmax}\left(\frac{\mathbf{Q}^{t}\left(\mathbf{K}^{t}\right)^{\mathrm{T}}}{\sqrt{d_{k}}}\right)\mathbf{V}^{t}$ (11)
-where $\mathbf{H}^{l}$ represents the input at the $l^{th}$ layer, $\mathbf{Q}^{t},\mathbf{K}^{t}$, and $\mathbf{V}^{t}$ are queries, keys, and values, respectively, derived from $\mathbf{H}^{l}$, and $d_{k}$ is the dimension of keys.
-The spatial attention mechanism focuses on modeling joint relationships in space:
-$\mathbf{Q}^{s},\mathbf{K}^{s},\mathbf{V}^{s}=\mathcal{F}^{s}_{QKV}\left(\mathbf{Z}^{t}\right)$ (12)
-$\mathbf{Z}^{s}=\operatorname{softmax}\left(\frac{\mathbf{Q}^{s}\left(\mathbf{K}^{s}\right)^{\mathrm{T}}}{\sqrt{d_{k}}}\right)\mathbf{V}^{s}$ (13)
-where $\mathbf{Q}^{s},\mathbf{K}^{s}$, and $\mathbf{V}^{s}$ are queries, keys, and values for spatial attention, respectively, similarly derived from $\mathbf{Z}^{t}$.
-$\mathbf{Z}^{s}$ is the output of ST-Attention.
-PMG-Attention
-PMG-Attention is employed within the Future Motion Predictor (FMP) module to incorporate past motion features into the prediction of future motion sequences.
-Unlike ST-Attention which focuses on the past motion sequence itself, PMG-Attention applies cross-attention between the current state and the past motion features:
-$\mathbf{Q}^{t},\mathbf{K}^{t},\mathbf{V}^{t}=\mathcal{F}^{t}_{QKV}\left(\mathbf{H}^{l}\right),\mathcal{F}^{t}_{QKV}\left(\mathbf{H}^{P}\right)$ (14)
-$\mathbf{Z}^{t}=\operatorname{softmax}\left(\frac{\mathbf{Q}^{t}\left(\mathbf{K}^{t}\right)^{\mathrm{T}}}{\sqrt{d_{k}}}\right)\mathbf{V}^{t}$ (15)
-$\mathbf{Q}^{s},\mathbf{K}^{s},\mathbf{V}^{s}=\mathcal{F}^{s}_{QKV}\left(\mathbf{Z}^{t}\right),\mathcal{F}^{s}_{QKV}\left(\mathbf{H}^{P}\right)$ (16)
-$\mathbf{Z}^{s}=\operatorname{softmax}\left(\frac{\mathbf{Q}^{s}\left(\mathbf{K}^{s}\right)^{\mathrm{T}}}{\sqrt{d_{k}}}\right)\mathbf{V}^{s}$ (17)
-where $\mathbf{Q}^{t}$ and $\mathbf{Q}^{s}$ are obtained from the hidden state of future motions, and the $\mathbf{K}^{t},\mathbf{V}^{t},\mathbf{K}^{s},\mathbf{V}^{s}$ are obtained from the features of past motion.
+Figure 7: Visualization of motion prediction results on the processed FineMotion dataset.
+Figure 8: Visualization of motion prediction results (“Posing”, “Walking”, “Greeting”) on Human3.6M dataset.
+Figure 9: Visualization of motion prediction results (“Posing”, “Walking”, “Greeting”) on Human3.6M dataset.
+In this supplementary material, we provide additional technical details, extended results, and further qualitative analyses that complement the main paper.
+Specifically, we include a detailed explanation of the spatiotemporal attention mechanisms used in our architecture (Sec.
+7), implementation details and training configurations (Sec.
+8), full motion prediction results on the Human3.6M dataset (Sec.
+13), comprehensive ablation studies on model components and architectural choices (Sec.
+10), and extensive visualizations, including motion prediction results, and attention feature maps (Sec.
+9 and Sec.
+12).
+We also discuss the mask strategy and reconstruction performance in Sec.
+11.
+Together, these supplementary materials offer a deeper understanding of our framework and provide further evidence of its effectiveness and robustness.
+7 Detailed Explanation of Model Architecture
+We provide additional details of the two attention modules used in our framework: the SpatioTemporal Attention (ST-Attention) in the Past Motion Encoder (PME), and the spatiotemporal cross-attention in the Future Motion Decoder (FMD).
+Both modules follow the standard Transformer design, where each attention sub-layer (temporal or spatial) is immediately followed by a position-wise feed-forward network (FFN), each equipped with its own residual connection and layer normalization.
+ST-Attention (within PME).
+Each ST-Attention layer contains a temporal attention block and a spatial attention block, applied sequentially.
+Given the input to the $l^{th}$ layer $\mathbf{H}^{l}$, the temporal self-attention is computed as:
+$\mathbf{Q}^{t},\mathbf{K}^{t},\mathbf{V}^{t}=\mathcal{F}^{t}_{QKV}(\mathbf{H}^{l}),$ (11)
+$\mathbf{A}^{t}=\operatorname{softmax}\!\left(\frac{\mathbf{Q}^{t}(\mathbf{K}^{t})^{\mathrm{T}}}{\sqrt{d_{k}}}\right)\mathbf{V}^{t}.$ (12)
+A position-wise FFN follows:
+$\mathbf{H}^{l,t}=\mathrm{LN}\!\left(\mathbf{A}^{t}+\mathrm{FFN}(\mathbf{A}^{t})\right).$ (13)
+The refined features $\mathbf{H}^{l,t}$ then pass through a spatial attention block:
+$\mathbf{Q}^{s},\mathbf{K}^{s},\mathbf{V}^{s}=\mathcal{F}^{s}_{QKV}(\mathbf{H}^{l,t}),$ (14)
+$\mathbf{A}^{s}=\operatorname{softmax}\!\left(\frac{\mathbf{Q}^{s}(\mathbf{K}^{s})^{\mathrm{T}}}{\sqrt{d_{k}}}\right)\mathbf{V}^{s},$ (15)
+followed again by an FFN:
+$\mathbf{H}^{l+1}=\mathrm{LN}\!\left(\mathbf{A}^{s}+\mathrm{FFN}(\mathbf{A}^{s})\right).$ (16)
+Spatiotemporal Cross-Attention (within FMD).
+FMD incorporates the encoded past motion features $\mathbf{H}^{P}$ via spatiotemporal cross-attention, where queries come from the current decoding state and keys/values come from the past sequence.
+As in PME, each temporal and spatial block has an independent FFN.
+Temporal cross-attention is given by:
+$\mathbf{Q}^{t}=\mathcal{F}^{t}_{Q}(\mathbf{H}^{l}),\qquad\mathbf{K}^{t},\mathbf{V}^{t}=\mathcal{F}^{t}_{KV}(\mathbf{H}^{P}),$ (17)
+$\mathbf{A}^{t}=\operatorname{softmax}\!\left(\frac{\mathbf{Q}^{t}(\mathbf{K}^{t})^{\mathrm{T}}}{\sqrt{d_{k}}}\right)\mathbf{V}^{t},$ (18)
+followed by:
+$\mathbf{H}^{l,t}=\mathrm{LN}\!\left(\mathbf{A}^{t}+\mathrm{FFN}(\mathbf{A}^{t})\right).$ (19)
+Spatial cross-attention further injects structural cues from past motion:
+$\mathbf{Q}^{s}=\mathcal{F}^{s}_{Q}(\mathbf{H}^{l,t}),\qquad\mathbf{K}^{s},\mathbf{V}^{s}=\mathcal{F}^{s}_{KV}(\mathbf{H}^{P}),$ (20)
+$\mathbf{A}^{s}=\operatorname{softmax}\!\left(\frac{\mathbf{Q}^{s}(\mathbf{K}^{s})^{\mathrm{T}}}{\sqrt{d_{k}}}\right)\mathbf{V}^{s},$ (21)
+with its FFN:
+$\mathbf{H}^{l+1}=\mathrm{LN}\!\left(\mathbf{A}^{s}+\mathrm{FFN}(\mathbf{A}^{s})\right).$ (22)
+This spatiotemporal cross-attention design enables FMD to effectively leverage both temporal evolution and spatial structure in the past sequence when generating future motions.
 8 Implementation Details
-We deployed our network using an NVIDIA RTX4090 GPU (with PyTorch 2.0).
-For training, we set the batch size to 24 and utilized the Adam optimizer with an initial learning rate of 0.0005.
-To enhance convergence, we implemented a cosine annealing learning rate schedule.
-These training parameters were consistently applied across all datasets in our experiments.
-We ensured that the GPU memory usage remained below 6GB, enabling our network to be seamlessly deployed on various hardware platforms with ease.
-9 Complete Results on Human3.6M
-We provide the full results of comparisons on the Human3.6M dataset in Tab. 7 as a supplement to Tab. 1. Despite the simplicity of our network structure, we achieved the best results on most actions.
+We implement all models using PyTorch 2.0 and train them on a single NVIDIA RTX4090 GPU.
+The batch size is set to 24, and the Adam optimizer is used with an initial learning rate of 0.0005.
+A cosine annealing schedule is adopted to gradually decay the learning rate, and the overall GPU memory consumption remains below 6GB, ensuring that our method can be easily deployed on a wide range of hardware platforms.
+For network configuration, the Past Motion Encoder (PME), Future Motion Decoder (FMD), and Future Semantic Unit (FSU) each contain 3 stacked modules.
+The feature dimension is kept constant at 128 within PME and FMD, and at 256 within FSU.
+We employ 8-head attention, with each head containing 32 channels.
+During pretraining, the reconstruction losses of past and future motions share equal weights (=1).
+In the finetuning stage, the loss weight of text generation is set to 0.1, and the parameters of the FSU are updated every 5 epochs.
+For self-supervised training on H36M, 3DPW, and AMASS, we train the model for 40 epochs and then finetune it for an additional 20 epochs on the motion prediction task.
+The best checkpoint on the validation set is selected after the self-supervised stage.
+For the processed FineMotion dataset, we train for 300 epochs in the self-supervised phase and subsequently finetune for 200 epochs on motion prediction.
+9 More visualization results
+We provide additional visualizations of motion prediction results on the processed FineMotion dataset in Fig.
+7, and further show qualitative results on the Human3.6M dataset in Fig.
+9 and Fig.
+8.
+On Human3.6M, our method accurately predicts joint movements and maintains consistent body configurations over time.
+In particular, fine-grained actions involving limb articulations, such as arm swings, remain temporally coherent and physically plausible.
+Compared with Human3.6M, the motions in FineMotion are more complex and the prediction horizon is significantly longer (up to 1.5 seconds).
+Even under these challenging conditions, our model successfully captures the correct motion tendencies and preserves the overall dynamics of large-amplitude arm movements and posture transitions.
+This robustness benefits from our self-supervised pretraining strategy, which encourages the model to learn implicit spatiotemporal constraints among joints.
+Nevertheless, we observe noticeable errors at distal joints, especially in highly articulated or fast-changing motion segments.
+This limitation partly arises from the inherent uncertainty and variability in the execution of fine-grained actions.
+In future work, we plan to address this issue by explicitly modeling motion uncertainty and incorporating stronger motion priors into the prediction framework.
 10 Ablation Study on Model Architecture
+Table 9: Ablation study on future text generation performance (future motion semantic understanding).
+Method BLEU@4 BLEU@1 ROUGE CIDEr
+w/o SSL 17.6 6.5 30.9 24.0
+w/ SSL 18.2 7.8 32.5 24.2
+Effect of SSL on Future Motion Understanding
+We conduct an ablation study to examine how our self-supervised training stage influences the model’s ability to understand future motion semantics.
+As shown in Tab.
+9, removing the SSL stage and directly training motion prediction and semantic understanding in a joint supervised manner (w/o SSL) leads to worse text generation performance.
+These results demonstrate that the SSL stage effectively enhances semantic reasoning about future dynamics and provides more informative motion representations for downstream captioning.
 Table 10: Ablation study of model architecture
@@
 We experimented with other fusion methods, with results shown in Tab.
-8: A) Direct addition of features from past and future actions; B) Concatenation of both in the feature dimension, followed by a Linear layer to reduce the number of channels back to the original value.
+10: A) Direct addition of features from past and future actions; B) Concatenation of both in the feature dimension, followed by a Linear layer to reduce the number of channels back to the original value.
 We observed that cross-attention is the best choice because it explicitly computes the temporal and spatial interactions between the past and future motion.
@@
 D complete 8.8 19.3 41.1 51.8 1.66M
-Figure 4: Visualization of attention feature maps.
 Spatiotemporal Attention
-In the default setting, we follow the AuxFormer[41] of computing temporal and spatial attention sequentially.
-Results of parallel computation of temporal and spatial attention are presented in Tab.
-8 C, showing worse performance.
+In our default configuration, we adopt the sequential computation scheme from AuxFormer [xu2023auxiliary], where temporal attention and spatial attention are applied one after the other to progressively capture motion dynamics and joint dependencies.
+In contrast, the parallel computation variant (see Tab.
+10 C), which processes temporal and spatial relations simultaneously, produces noticeably inferior performance, indicating that the sequential formulation provides a more effective inductive bias for structured human motion modeling.
 Model Size
 In Tab.
-9, we show the impact of different model sizes (including feature dimension $d$ in attention layer and number of layers $l$) on parameter count and results.
+11, we show the impact of different model sizes (including feature dimension $d$ in attention layer and number of layers $l$) on parameter count and results.
 Although reducing the model size can decrease the number of parameters, it can also lead to decreased performance.
-We observe that the default setting ($d$=256, $l$=3) represents the optimal balance between parameter count and performance.
+We observe that the default setting (dd=256, ll=3) represents the optimal balance between parameter count and performance.
+Figure 10: Visualization of global per-joint masking probability and the corresponding reconstruction error.
+Figure 11: Visualization of reconstruction results. The blue lines are ground-truth value and the red lines represent the reconstruction motion skeleton sequences.
+Table 12: Per-action reconstruction MPJPE at different time horizons ($m$).
+Action 80ms 160ms 320ms 400ms
+walking 2.55 6.25 12.54 15.48
+eating 1.51 3.47 7.04 8.81
+smoking 1.57 3.42 6.17 7.65
+discussion 1.90 4.33 7.66 9.54
+directions 1.45 3.41 7.39 9.45
+greeting 2.45 6.08 12.52 15.64
+phoning 1.88 4.47 9.28 11.63
+posing 2.18 4.99 8.84 11.12
+purchases 2.31 5.44 11.37 14.80
+sitting 2.02 4.65 9.63 12.30
+sittingdown 2.50 4.86 7.91 9.83
+takingphoto 1.81 4.17 8.93 11.59
+waiting 1.92 4.52 9.19 11.7
+walkingdog 4.16 10.37 19.73 24.23
+walkingtogether 1.91 4.89 10.92 14.00
+Average 2.14 5.02 9.94 12.82
+11 Analysis of Velocity-Based Joint Masking and Reconstruction Performance
+To better understand how velocity-driven masking influences representation learning, we compute the global per-joint masking probability and the corresponding reconstruction error across the entire Human3.6M test set.
+The results are shown in Fig.
+10.
+First, the masking distribution (left) exhibits a clear semantic pattern: High-velocity distal joints such as the feet, wrists, hands, and fingers are assigned higher masking probabilities; low-motion central joints, such as the spine, thorax, and neck, receive much smaller masking rates.
+This confirms that our velocity-based strategy automatically focuses learning pressure on dynamically informative regions of the body, rather than applying a uniform mask.
+Such non-uniform perturbation encourages the model to learn motion-consistent representations that reflect true physical dynamics, instead of memorizing static structure.
+Second, the reconstruction error plot (right) shows a correlation with the masking distribution.
+Highly dynamic joints not only move faster but also demonstrate larger pose variability, making them inherently more challenging to reconstruct.
+Interestingly, although these dynamic joints are frequently masked, their reconstruction error does not explode.
+As shown in the right figure, the model maintains a stable reconstruction range (approximate 6 mm to 12 mm).
+This is largely because their adjacent, low-velocity joints remain visible and provide strong spatial constraints, allowing the model to leverage local kinematic continuity for recovery.
+In other words, the structured skeletal graph prevents error propagation and enables the model to infer masked high-frequency motions using nearby static or slow-moving joints.
+The qualitative reconstructions in Fig.
+11 further validate the effectiveness of our masking strategy.
+Even when large portions of the distal limbs are removed, the model is able to faithfully recover the full-body motion trajectory, producing temporally coherent poses that closely match the unmasked ground truth.
+Notably, complex motion cues—such as arm swings in Posing, subtle conversational gestures in Discussion, and periodic gait cycles in Walking—are accurately restored, demonstrating the model’s ability to infer fine-grained dynamics from limited visible context.
+These visual results highlight the robustness of our learned representations and showcase how motion structure priors guide reliable completion of challenging high-frequency movements.
+The per-action results in Tab.
+12 show that dynamic actions with large limb motions incur higher reconstruction errors, while static actions remain easy to recover.
+Despite this variation, our velocity-aware masking maintains stable performance across both low- and high-motion activities.
+Figure 12: Visualization of attention feature maps.
 12 Visualization of Attention Feature Map
@@
 As shown in Fig.
-4, sub figure (a) and (b) display the temporal self-attention and spatial self-attention feature maps from different layers in PME, sub figure (c) and (d) show the temporal cross-attention and spatial cross-attention feature maps from different layers in FMP.
+12, sub figure (a) and (b) display the temporal self-attention and spatial self-attention feature maps from different layers in PME, sub figure (c) and (d) show the temporal cross-attention and spatial cross-attention feature maps from different layers in FMP.
 In temporal cross-attention, the model’s first and third layers tend to allocate attention to the later frames, as these frames have a greater influence on future sequences.
@@
 In spatial cross-attention, the model identifies significant correlations between the movement patterns of the ’thorax’, ’head’, ’left elbow’, and ’right elbow’ joints and the prediction of future actions.
+13 Full Results on Human3.6M
+We provide the full results of comparisons on the Human3.6M dataset in Tab. 13 as a supplement to Tab. 1. Despite the simplicity of our network structure, we achieved the best results on most actions.
+Table 13: Comparisons of MPJPEs between our proposed method with other state-of-the-art methods for short-term prediction on Human3.6.
+We show the results across all actions such as walking and eating.
+The best results are highlighted in bold, and the second best results is underlined
+Motion Walking Eating Smoking Discussion Direction
+Milliseconds 80 160 320 400 80 160 320 400 80 160 320 400 80 160 320 400 80 160 320 400
+Mixer[bouazizi2022motionmixer] 12.0 23.0 42.4 51.3 7.8 16.5 34.3 43.0 7.4 15.2 30.3 37.6 10.7 25.0 56.9 71.1 7.5 18.4 43.9 55.3
+PGBIG[ma2022progressively] 10.2 19.8 34.5 40.3 7.0 15.1 30.6 38.1 6.6 14.1 28.2 34.7 10.0 23.8 53.6 66.7 7.2 17.6 40.9 51.5
+SPGSN[li2022skeleton] 10.1 19.4 34.8 41.5 7.1 14.8 30.5 37.9 6.7 13.8 28.0 34.6 10.4 23.8 53.6 67.1 7.4 17.2 40.0 50.3
+siMLPe[guo2023back] 11.9 19.3 31.9 36.9 10.2 16.5 30.7 37.3 10.4 16.4 29.7 35.9 12.7 25.3 55.1 68.1 11.0 19.4 41.6 52.1
+DSTD-GC[fu2023learning] 11.1 22.4 38.8 45.2 7.0 15.5 31.7 39.2 6.6 14.8 29.8 36.7 10.0 24.4 54.5 67.4 6.9 17.4 40.1 51.7
+AuxFormer[xu2023auxiliary] 8.9 16.9 30.1 36.1 6.4 14.0 28.8 35.9 5.7 11.4 22.1 27.9 8.6 18.8 38.8 49.2 6.8 17.0 40.3 51.6
+CIST-GCN[medina2024context] 11.8 23.4 40.5 46.5 6.7 14.8 29.8 36.8 7.3 15.6 31.0 38.0 10.2 23.7 52.3 65.3 7.3 18.1 43.6 55.3
+GCNext[wang2024gcnext] 8.8 38.9 5.9 35.0 5.6 36.1 8.8 63.1
+ALIEN[wei2025alien] 9.5 18.7 33.8 40.2 6.6 14.6 30.5 38.1 6.4 13.9 28.8 35.7 9.6 23.7 53.8 67.2 6.6 16.8 40.5 51.4
+Ours 8.6 16.5 31.2 37.7 6.1 13.7 29.3 36.8 5.3 10.3 19.7 25.2 7.7 16.3 31.7 40.9 6.3 15.8 39.8 51.4
+Motion Greeting Phoning Posing Purchases Sitting
+Milliseconds 80 160 320 400 80 160 320 400 80 160 320 400 80 160 320 400 80 160 320 400
+Mixer[bouazizi2022motionmixer] 16.4 37.1 80.2 98.2 9.1 19.5 42.4 53.6 11.5 27.0 64.9 83.5 13.4 30.7 65.6 80.0 9.8 20.9 45.6 57.3
+PGBIG[ma2022progressively] 15.2 34.1 71.6 87.1 8.3 18.3 38.7 48.4 10.7 25.7 60.0 76.6 12.5 28.7 60.1 73.3 8.8 19.2 42.4 53.8
+SPGSN[li2022skeleton] 14.6 32.6 70.6 86.4 8.7 18.3 38.7 48.5 10.7 25.3 59.9 76.5 12.7 28.6 61.0 74.4 9.3 19.4 42.2 53.6
+siMLPe[guo2023back] 16.2 33.7 70.9 86.3 11.9 20.3 39.7 48.9 13.9 27.3 62.0 78.6 15.5 29.3 59.3 72.0 13.2 21.9 44.5 55.7
+DSTD-GC[fu2023learning] 14.3 33.5 72.2 87.3 8.5 19.2 40.3 49.9 10.1 25.4 60.6 77.3 12.7 29.7 62.3 75.8 8.8 19.3 42.9 54.3
+AuxFormer[xu2023auxiliary] 13.5 31.3 69.2 85.4 7.9 17.3 37.4 47.2 8.8 19.1 39.2 51.0 11.9 28.0 61.8 76.3 8.7 19.0 42.1 53.3
+CIST-GCN[medina2024context] 13.7 31.0 65.7 79.9 8.6 18.5 39.3 49.6 9.6 23.7 57.7 75.0 13.3 30.2 63.0 77.3 8.9 19.4 42.3 53.6
+ALIEN[wei2025alien] 13.6 32.1 70.3 85.8 8.1 18.2 39.3 49.4 9.8 25.1 61.2 78.9 12.2 28.9 62.0 76.6 8.7 19.2 42.9 54.9
+Ours 10.8 26.4 62.7 79.8 7.5 17.1 37.9 48.0 8.2 16.9 32.8 43.5 11.4 27.6 62.6 77.6 8.5 19.2 44.3 56.7
+Motion Sitting Down Taking Photo Waiting Walking Dog Walking Together
+Milliseconds 80 160 320 400 80 160 320 400 80 160 320 400 80 160 320 400 80 160 320 400
+Mixer[bouazizi2022motionmixer] 15.3 30.5 62.7 78.0 9.0 19.9 44.7 56.9 9.9 22.0 49.7 62.5 21.1 43.9 84.7 101.2 9.9 20.4 39.3 47.3
+PGBIG[ma2022progressively] 13.9 27.9 57.4 71.5 8.4 18.9 42.0 53.3 8.9 20.1 43.6 54.3 18.8 39.3 73.7 86.4 8.7 18.6 34.4 41.0
+SPGSN[li2022skeleton] 14.2 27.7 56.7 70.7 8.8 18.9 41.5 52.7 9.2 19.8 43.1 54.1 8.9 18.2 33.8 40.9
+siMLPe[guo2023back] 18.0 29.8 57.7 71.2 13.1 21.3 42.9 53.6 11.9 21.5 44.4 55.0 21.8 40.6 74.6 87.0 11.9 19.8 34.2 39.8
+DSTD-GC[fu2023learning] 14.1 28.0 57.3 71.2 8.4 18.8 42.0 53.5 8.7 20.2 44.3 55.3 19.6 41.8 77.6 90.2 9.1 19.8 36.3 42.7
+AuxFormer[xu2023auxiliary] 13.5 27.6 57.7 72.2 8.2 18.4 41.5 53.0 8.2 18.5 41.2 52.2 17.1 36.5 70.4 83.0 7.8 15.9 30.2 37.0
+CIST-GCN[medina2024context] 14.1 29.8 57.3 69.8 8.2 18.4 40.6 51.8 8.6 19.4 43.5 54.8 20.0 41.4 73.7 85.1 9.6 20.3 38.2 45.6
+ALIEN[wei2025alien] 13.4 27.0 56.4 70.5 8.8 19.0 42.1 53.5 8.3 19.2 43.3 54.8 18.7 40.3 75.9 88.5 8.4 17.8 33.7 40.4
+Ours 10.1 17.2 30.2 39.2 7.6 17.8 42.0 54.1 7.8 18.0 41.6 53.2 19.1 41.1 80.6 95.2 7.3 15.3 30.8 37.9